### v2

# PIMRL: Physics-Informed Multi-Scale Recurrent Learning for Spatiotemporal Prediction

Abstract
Simulation of spatiotemporal systems governed by partial differential equations is widely applied in fields such as biology, chemistry, aerospace dynamics, and meteorology. Traditional numerical methods incur high computational costs due to the requirement of small time steps for accurate predictions. While machine learning has reduced these costs, long-term predictions remain challenged by error accumulation, particularly in scenarios with insufficient data or varying time scales, where stability and accuracy are compromised. Existing methods often neglect the effective utilization of multi-scale data, leading to suboptimal robustness in predictions. To address these issues, we propose a novel multi-scale learning framework, namely, the Physics-Informed Multi-Scale Recurrent Learning (PIMRL), to effectively leverage multi-scale data for spatiotemporal dynamics prediction. The PIMRL framework comprises two modules: the micro-scale module embeds physical knowledge into neural networks via pretraining, and the macro-scale module adopts a data-driven approach to learn the temporal evolution of physics in the latent space. Experimental results demonstrate that the PIMRL framework consistently achieves state-of-the-art performance across five benchmark datasets ranging from one to three dimensions, showing average improvements of over 9% in both RMSE and MAE evaluation metrics, with maximum enhancements reaching up to 80%.
1 Introduction
In the field of natural sciences, physical systems governed by partial differential equations (PDEs) have found widespread applications across disciplines including biology, chemistry, meteorology, etc. [1, 2, 3, 4, 5]. Although numerical methods have been regarded as reliable tools for modeling these systems, the use of Direct Numerical Simulation (DNS) faces significant hurdles due to inherent limitations. DNS necessitates high spatial resolution and fine time stepping, resulting in considerable computational demands and prolonged processing times. A case in point is the simulation of aerodynamic flows around aircraft, which typically requires the generation of millions of grid points, thereby imposing prohibitive computational requirements [6, 7]. Moreover, these numerical simulation methods require complete physical prior knowledge, such as PDE formula, parameters, and initial/boundary conditions [8].
The ongoing development of artificial intelligence (AI) has propelled research into data-driven simulation methods, showcasing significant potential [9, 10, 11, 12]. These methods, which do not require prior physics knowledge, offer user-friendly solutions. They also overcome traditional constraints on resolution and small time stepping, ensuring accurate solutions. Nevertheless, when faced with sparse data and multi-scale temporal challenges, these methods often struggle to optimally utilize data. They may either discard micro-scale data in favor of macro-scale data or vice versa, leading to compromised accuracy.
Multi-scale Burst sampling is a technique capturing multiple samples at a high rate over a short period of time to record rapidly changing events or transient phenomena, e.g., fast dynamics, followed by a low sampling rate to capture slow dynamics (see Figure 1). Compared to traditional continuous sampling methods, it provides high-resolution data at critical moments while maintaining resource efficiency. But conventional uniform-scale models struggle to fully utilize such data. The latent ODE approach [13] may work for multi-scale sampled data, but remains unclear whether it can address PDE problems. There is an urgent need for a new method to handle such a type of data for spatiotemporal systems.
Figure 1: Multi-scale sampling, where $\Delta\tau$ denotes the micro-scale time interval for fast dynamics, $\Delta t$ the macro-scale time interval for slow dynamics, and $\zeta_{t}$ the scale separation variable (typically $\zeta_{t}<1$ or $\zeta_{t}\ll 1$).
Data-driven methods face significant limitations when addressing challenges related to insufficient data and multiple time scales mentioned above. To overcome these obstacles, current approaches integrate physical knowledge into the model learning process. Specifically, physics-informed neural networks (PINNs) [14, 15, 16, 17, 18] design initial and boundary conditions as penalty terms in the loss function, thereby leveraging physical laws in a “soft” way. However, this soft embedding approach can sometimes lead to unsatisfactory results. Also, to some extent, the embedding methods used in PINNs are unable to ensure that the model fully adheres to the embedded physical conditions.
Directly embedding physical equations into the model architecture, as seen in the physics-encoded recurrent convolutional neural network (PeRCNN) [19], ensures strict adherence to the underlying physical laws. This approach addresses the issue of soft constraints in PINNs and enhances the model’s interpretability and generalization capabilities. However, this approach requires continuous micro-scale data, which imposes stringent demands on the dataset and can introduce instability in long-term predictions. Alternatively, methods such as the learned interpolation (LI) model [20] and multiscale simulation frameworks [21] integrate numerical techniques with neural networks. These hybrid approaches enhance simulation efficiency while maintaining acceptable accuracy, although they still have some of the limitations inherent in traditional numerical methods.
Moreover, we observe that the error increases along with the prediction horizon, a phenomenon that is widespread. Although existing efforts are predominantly placed on enhancing the model’s predictive capability, long-term prediction of spatiotemporal dynamics still suffers from error accumulation. Hence, properly controlling the accumulation of errors is becomes crucial. Based on this core idea, we are motivated to leverage comprehensive multi-scale data through a novel Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework to tackle the problems mentioned above. This framework consists of macro-scale and micro-scale modules. The micro-scale module, pretrained to learn the underlying physical laws, enhances the accuracy of simulations. The macro-scale module reduces the accumulation of errors by minimizing the number of rollout iterations for the micro-scale module, thereby enhancing long-term predictive performance. The main contributions of this paper are summarized as follows:
  • We proposed a new PIMRL framework that effectively leverages information from multi-scale data for long-term spatiotemporal dynamics prediction. The concept of reducing the accumulation of errors is achieved through the integration of the macro-scale and micro-scale modules.
  • We designed a novel message passing mechanism between micro- and macro-scale modules, which effectively transmits physical information, enhances the micro-module’s correction capability, and reduces the number of micro-scale rollout iterations through the macro-scale module.
  • The PIMRL model achieved optimal performance in effectively predicting the duration of multiple different cases from fluid dynamics to physical systems, demonstrating its scalability and laying a solid foundation for more generalizable models in the future.
2 Related Work
Simulation tasks often aim to solve partial differential equations (PDEs) accurately and efficiently. Previous researchers have developed numerical methods achieving high precision with many nodes and short time steps. To speed up simulations, deep learning techniques have evolved from applying conventional algorithms to designing new models that integrate physical knowledge, including hybrid methods and models embedding physical principles.
Computational fluid dynamics. Computational Fluid Dynamics (CFD) is a branch of fluid mechanics that uses numerical methods and algorithms to analyze and solve problems involving fluid flow [1, 2, 3, 4, 5]. When faced with complex problems, unacceptable time and computational costs are the primary obstacles limiting the application of CFD. Moreover, when the corresponding physical equations contain unknown parameters or even unknown terms, numerical methods are unable to perform accurate simulations. The aforementioned methods have inspired us to leverage physical knowledge. However, to overcome the aforementioned issues, we have decided to introduce deep learning methods into our PIMRL framework.
Deep learning methods. With the advancement of AI, the application of AI in physical system simulation tasks has become more diverse and profound. For example, classical convolutional neural network models [11, 22] and ResNet [23], U-Net [24] what originally used for image segmentation, graph neural networks [25, 26], and transformer-based models [27, 28, 29, 30] are now being employed. Neural operators used for learning mappings between function spaces have also seen significant development in physical simulation tasks, such as DeepONet [9], MWT [31], FNO [10, 32, 33, 34], etc. In addition, there are several methods specifically designed for spatiotemporal prediction tasks, such as ConvLSTM [35], PredRNN [36], and TrajGRU [37]. The traditional deep learning methods mentioned above all require sufficient data for training; insufficient data can lead to poor model performance and severe overfitting issues. Moreover, error accumulation is another inevitable problem that limits the application of traditional deep learning methods. These limitations restrict the effectiveness and reliability of such models, especially in scenarios requiring long-term predictions or when data is scarce. The challenges outlined above have led us to propose a framework that effectively controls error accumulation and adeptly processes multi-scale data.
Physics-informed deep learning methods. To incorporate physical information into models, researchers have devised various methods. One category includes physics-inspired methods like PhyCRNet [38], PINN [14, 15, 16, 17, 18], PhyLSTM [39], PhySR [40], and PhyMPGN [41]. Another category involves physical embedding methods, where explicit physics knowledge is embedded into the model to fully leverage physical principles, like EquNN [42] and PDE-Net [43, 44]. PeRCNN [19, 45] offers a “hard” encoding mechanism to learn the dynamics of physical systems from limited data. This approach leverages prior physics knowledge for predictions, thereby equipping strong predictive capabilities and robust generalization across different initial conditions. However, all the aforementioned methods suffer from error accumulation, making it difficult to obtain stable and accurate results in long-term prediction tasks. PeRCNN effectively utilizes physical information in a way that can be leveraged within PIMRL.
Hybrid learning methods. In recent years, an emerging research direction has been to integrate deep learning methods with traditional numerical methods. By combining classical solvers with deep neural networks, hybrid approaches designed on this principle can achieve acceptable accuracy while operating faster than pure numerical methods. Examples include the Learned Interpolation (LI) model [20], numerical discretization learning [46], and P2C2Net [47]. However, in these hybrid methods, the numerical method component is not involved in the training process, and such approaches also require substantial amounts of data. The pioneering work multi-scale simulations of complex systems [21] employs a multi-scale framework for predictions, effectively reducing the required time and computational cost while maintaining good accuracy. However, this method still requires ample data for training. The idea behind hybrid learning methods is very promising, so we have introduced our own message-passing mechanism within our PIMRL framework.
3 Methodology
We propose the PIMRL framework, as illustrated in Figure 2, for spatiotemporal dynamics prediction with a small amount of multi-scale training data
3.1 Partial Differential Equations simulation task
Simulation tasks are intimately connected with PDEs, which are essential for describing and simulating physical models, and the time-dependent PDEs are defined as:
$\mathbf{u}_{t}=\mathcal{F}(t,x,\mathbf{u},\nabla\mathbf{u},\mathbf{u}\cdot\nabla\mathbf{u},\nabla^{2}\mathbf{u},\cdot\cdot\cdot;\mu),$ (1)
where $\mathbf{u}(x,t)$ denotes the spatiotemporal solution field, $\mathbf{u}_{t}$ the first-order time derivative, $\mathcal{F}(\cdot)$ a linear/nonlinear function, $\nabla$ the Nabla operator, $\nabla^{2}$ the Laplace operator, and $\mu$ the PDE parameter.
In addition, we define the initial and boundary conditions (ICs, BCs) for this equation, namely: $I[\mathbf{u}](x,t=0)=0$ and $B[\mathbf{u}](x,t)=0$ where $I$ and $B$ indicate IC and BC operators.
3.2 Overview
In this study, we aim establish a model capable of predicting the evolution of spatiotemporal dynamics for nonlinear PDE systems at a long-term horizon, given limited training data (e.g., a few trajectories resulted from different initial conditions). As outlined in the introduction, we have designed the PIMRL framework to mitigate error accumulation while ensuring that it can effectively capture the intrinsic changes in the physical system, rather than disregarding the physical information over large time intervals. Additionally, PIMRL is capable of efficiently utilizing multi-scale data. Below, we will provide a detailed exposition of the PIMRL framework’s architecture, the corresponding training methodologies, the macro and micro modules, as well as the boundary padding method that incorporates boundary conditions.
Figure 2: The overall framework architecture, which integrates physics-informed constraints with deep learning.The initial state of the system is denoted by $\mathbf{u}_{0}$. The state predicted by the micro-module after $k$ iterations, where each iteration occurs at intervals of $\delta t$, is represented by $\mathbf{u}^{micro}_{k\delta t}$. The predicted value of the physical state from PIMRL is denoted by $\hat{\mathbf{u}}$.
3.3 Forecasting Architecture
As shown in Figure 2, the PIMRL framework comprises two key components: the micro-scale module and the macro-scale module. The message-passing mechanism is an interaction between the micro-module and the macro-module. The information transmitted includes physical knowledge learned by the micro-module and corrections applied to the macro-module. Additionally, the macro-module also passes information to future iterations of the micro-module.
PIMRL is designed to achieve long-term prediction of spatiotemporal dynamics. PIMRL operates in a recursive manner, involving rollouts at the micro-module level, the macro-module level, and the overall framework level. Only the output from the macro-module contributes to the final output of the PIMRL framework and is used to compute the loss for training the entire framework, whereas the output from the micro-module serves as intermediate variables within the framework. And the message-passing mechanism is described as follows:
  • Firstly, the micro-module rollout is a simple autoregressive process with time step $\delta t$, where the output at the previous time step serves as the input for the next time step.
  • When the micro-module is involved in the prediction, for every $k$ steps of micro-module with $\delta t$ like Equation 3, the final output of micro-module is passed to the macro-module, and at this point, the output from the macro-module serves as the output of the entire PIMRL model shown as Equation 4. When the micro-module is not involved in the prediction, the macro-module rollout is a simple autoregressive process with time step $\Delta t$.
  • Finally, there is a PIMRL loop that operates in conjunction with the macro-module rollout. After every $N-1$ rollout of the macro-module, the micro-module stops participating in the prediction, and the macro-module performs $N$ steps of autoregressive prediction on its own. This completes a total of $2N$ rollouts. Each output from the macro-module during these $2N$ rollouts serves as the output of PIMRL as depicted in Equation 2.
We utilize two equations to represent the relationship between the micro-scale module and the macro-scale module, denoted as $F_{micro}$ and $F_{macro}$ respectively, as illustrated in Figure 2. The details of this process can be represented using the aforementioned symbols as follows:
[…]
where the variable $\mathbf{u}$ denotes the physical state. The relationship between the micro-time step $\delta t$ and the macro-time step $\Delta t$ is given by $\Delta t=k\delta t$, where $k$ is an adjustable parameter determined by the time stepping of different scales in the real data.
3.4 Training Strategies
Since the PIMRL model consists of micro- and macro-scale modules, adopting brute-force end-to-end training yields unsatisfactory results (see the ablation study). Firstly, we establish a pre-training phase where only the micro-module is trained. The purpose of this pre-training phase is to enable the micro-module to effectively learn the dynamics of the physical system and the underlying physical laws, free from the influence of the macro-module. In the next phase, referred to as the overall training phase, all modules within the PIMRL framework are engaged in training together. During this phase, the micro-module benefits from the parameters pre-trained in the pre-training phase, serving to supervise and correct the macro-module. The output from the macro-module serves as the output of the entire PIMRL for computing the loss. The details are shown in Appendix.
Figure 3: PIMRL includes two main modules: (a) the micro-module, designed to capture local features and small-scale dynamics; and (b) the macro-module, which captures long-term dependencies and global patterns using residual connections.
3.5 Micro-scale Module
The micro-scale module is designed to learn underlying physical laws that govern the spatiotemporal dynamics from micro-scale data with small time stepping, where we adopt the PeRCNN model [19] with the architecture of $\Pi$-block shown in Figure 3(a). In a forward Euler scheme: $\mathbf{u}_{(k+1)\delta t}=\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})\cdot\delta t+\mathbf{u}_{k\delta t},$ where $\delta t$ denotes that the module predicting in micro-scale time stepping. We can then approximate the $\mathcal{F}$ by $\hat{\mathcal{F}}$ described as follows:
^$\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})=\sum^{N_{c}}_{c=1}W_{c}\cdot\left[\prod_{N_{l}}^{l=1}(K_{c,l}\star\hat{\mathbf{u}}_{k\delta t}+b_{l})\right].$ (5)
where $N_{c}$ denotes the channel count, and $N_{l}$ the total number of parallel convolutional layers. The symbol $\star$ denotes the convolutional operation. For each layer $l$ and channel $c$, $K_{c,l}$ designates the specific filter weight, while $b_{l}$ stands for the bias term of that layer $l$. In the context of a $1\times 1$ convolutional layer, $W_{c}$ denotes the weight assigned to the $c^{\text{th}}$ channel, with the bias term being omitted here for the sake of simplicity and brevity.
When a certain term in the governing PDE remains known (e.g., the diffusion term $\Delta\mathbf{u}$), its discretization can be directly embedded in PeRCNN (called the physics-based Conv layer as shown in Figure 3(a)). The convolutional kernel in such a layer can be set according to the corresponding finite difference (FD) stencil. In essence, the physics-based Conv connection is constructed to incorporate known physical principles, whereas the $\Pi$-block is aimed at capturing the complementary unknown dynamics. The details of the physics-based FD Conv are provided in Appendix.
3.6 Macro-scale Module
The design of the macro-scale module, as a pivotal component of the PIMRL framework, is meticulously crafted to effectively manage and analyze macro-scale data. This type of data often poses unique challenges due to the substantial time spans it encompasses, which in turn lead to significant variations in the underlying physical states captured within the data points. These variations might be highly nonlinear and dynamic, making it difficult for traditional, physically motivated modeling methods to accurately capture all the nuances and complexities involved. As depicted in Figure 3, our macro-scale module utilizes ConvLSTM block with a residual connection. The structure of each block, illustrated in Figure 3(b), consists of a pair of encoders and decoders, along with a ConvLSTM cell shown in Appendix Figure. The input to this block undergoes mapping by the encoder, where the feature map serves as the characteristic of the latent space. Following this, the ConvLSTM cell simulates the dynamics, and the output is mapped back to the physical space through the decoder. Finally, the output, as a residual, is added to the feature from the previous time step to generate the prediction for the next time step. More details are provided in Appendix.
Figure 4: Periodic BC padding.
3.7 Boundary Condition Padding
Inspired by PeRCNN [19], we introduce BC hard encoding in both micro-scale and macro-scale modules. This encoding method ensures that the feature maps comply with the given BCs during the convolution process while also serving the purpose of padding, which involves filling the feature maps before convolution operations. Specifically, in this paper, our case adheres to periodic BCs, and the application of the corresponding padding is illustrated in Figure 4. This encoding scheme ingeniously incorporates BCs into the padding, thereby enhancing the accuracy of the prediction.
4 Experiments
To validate the effectiveness and versatility of our proposed PIMRL framework, we conducted extensive experiments on a diverse set of fluid dynamics and reaction-diffusion systems equations. Specifically, we tested our model on the following cases: the 1D Korteweg-de Vries (KdV) equation, the 2D Burgers equation, and three reaction-diffusion (RD) equations. These equations represent a range of physical phenomena with varying degrees of complexity and nonlinearity. Our results show that PIMRL consistently outperforms existing methods in terms of accuracy and robustness across these challenging cases.
Datasets. We conducted experiments on five datasets, including: Korteweg-de Vries (KdV), 2D Burgers, FitzHugh-Nagumo (FN), 2D Gray-Scott (2D GS), and 3D Gray-Scott (3D GS). The IC for the Kdv equation is created by summing multiple sine waves with random amplitudes, phases, and frequencies, resulting in a complex waveform. ICs for the Burgers’ equation are generated randomly according to a Gaussian distribution. The FN equation is initialized with random Gaussian noise for a warm-up period, after which time sequences are extracted to form the dataset. The GS equation starts the reaction from random initial positions and then diffuses. In these cases, except for the Kdv case, which was solved using the Finite Volume Method (FVM), the rest were solved using the Finite Difference Method (FDM). Additionally, We have two sets of data with different time scales originating from the same ICs. The micro-scale data $\textbf{U}^{\text{micro}}=\{\textbf{u}_{0},\textbf{u}_{\delta t},\textbf{u}_{2\delta t}\cdot\cdot\cdot\cdot\}\in\text{R}^{\text{micro}}$ is characterized by short and scattered continuous time intervals, while the macro-scale data $\textbf{U}^{\text{macro}}=\{\textbf{u}_{0},\textbf{u}_{\Delta t},\textbf{u}_{2\Delta t}\cdot\cdot\cdot\cdot\textbf{u}_{\text{Tend}}\}\in\text{R}^{\text{macro}}$ exhibits persistent continuity until the end. The validation and test sets are established based on different ICs but with the same parameters, making it more challenging than extrapolation under the same ICs. The dataset configuration is presented in Appendix.
Figure 5: An overview of the comparison between our PIMRL framework and baseline models is provided, including error propagation curves (left), error boxplots (middle), and final prediction plots (right). Figures (a) through (e) respectively showcase the qualitative results for the KdV, Burgers, 2D GS, FN, and 3D GS cases.
Model training. The primary objective is to first pretrain the micro-scale module using the micro-scale data to learn the underlying physics. Subsequently, the pretrained micro-scale module is integrated into the overall framework, and the entire model is trained using macro-scale data to capture the spatial evolution patterns over long time stepping. Both the pretraining and training processes can be formulated as auto-regressive rollout problems. The loss function is defined as $L(\boldsymbol{\theta})=\frac{1}{BN}\sum_{i=1}^{B}\sum_{j=1}^{N}(y_{i,j}-\hat{y}_{i,j})^{2}$, where the $B$ and $N$ denote the number of batches and the batch size. The $\boldsymbol{\theta}$ indicates all the trainable parameters and $(y_{i,j}-\hat{y}_{i,j})$ means the difference between the rollout-prediction $\hat{y}_{i,j}$ of $j$-th sample in the $i$-th batch and the corresponding labeled data $y_{i,j}$. More training details are shown in Appendix.
Baseline models. To validate the effectiveness of the proposed PIMRL framework, we introduced several baseline models. Firstly, we considered the widely recognized high-performing data-driven model FNO [10], which has been trained on datasets with two different time intervals, denoted as FNO (trained on fine-scale data with small time steps) and FNO-coarse (trained on coarse-scale data with large time steps). In particular, we examined the impact of cumulative error on long-term predictions. Our results show that the influence of error accumulation is significant. Detailed analysis is provided in Section 4.1. Secondly, we included the PeRCNN model [19], which embeds physical knowledge in a hard way and demonstrates excellent performance on multiple datasets. Due to constraints on the time stepping of the model, PeRCNN is trained only on datasets with small time stepping. Lastly, we incorporated ConvLSTM [35], a classic sequential prediction model trained on datasets with large time stepping, serving as the macro-scale module within our framework in a residual way. The details are shown in the Appendix.
Evaluation Metrics. To comprehensively evaluate the performance of our model, we adopted several metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and High Correction Time (HCT). These metrics provide a multi-faceted assessment of the model’s accuracy and reliability. RMSE is calculated on the macro-scale data to facilitate comparisons between models operating at different granularities. MAE provides a measure of the average absolute difference between the predicted and actual values,less sensitive to outliers compared to RMSE. HCT evaluates the time it takes for the model to correct its predictions to a high level of accuracy, which is particularly important for long-term prediction. Detailed formulas for these metrics are provided in Appendix.
4.1 Main Results
Figure 5 illustrates the performance results of our framework in comparison with various baseline models across different cases, including error propagation, showcasing box plots of different quantities and their prediction outcomes. Additionally, Table 1 presents relevant quantitative metrics as an expression of the results, where * in Table 1 indicates that the inference process has reached the end of the test data.
Table 1: Quantitative results of our model and baselines.
Case Model RMSE $\downarrow$ MAE $\downarrow$ HCT (s) $\uparrow$
KdV FNO 0.4891 0.3300 0.45
FNO-coarse 0.5461 0.4167 7.8
PeRCNN 0.0942 0.0941 30
PIMRL(Ours) $\mathbf{0.0457}$ $\mathbf{0.0607}$ $\mathbf{46.2}$
Promotion ($\uparrow$) 51.5% 35.5% 54.0%
Burgers FNO 0.1561 0.1301 0.104
FNO-coarse 0.1094 0.0879 0.064
PeRCNN 0.0075 0.0058 3.216
PIMRL(Ours) $\mathbf{0.0068}$ $\mathbf{0.0049}$ $\mathbf{3.216}$
Promotion ($\uparrow$) 9.3% 15.6% 0%
FN FNO 937 2393980 1.65
FNO-coarse 0.1878 0.1643 4.98
PeRCNN 0.1591 0.1139 6.99
PIMRL(Ours) $\mathbf{0.1349}$ $\mathbf{0.0990}$ $\mathbf{7.74}$
Promotion ($\uparrow$) 15.2% 13.1% 10.7%
2DGS FNO NaN NaN 810
FNO-coarse 0.0884 0.0629 1335
PeRCNN 0.0455 0.0268 1379.5
PIMRL(Ours) $\mathbf{0.0133}$ $\mathbf{0.0072}$ $\mathbf{1965^{*}}$
Promotion ($\uparrow$) 70.8% 73.1% 42.5%
3DGS FNO 0.2798 0.1950 112.5
FNO-coarse 0.1042 0.0611 360
PeRCNN 0.0532 0.0977 510
PIMRL(Ours) $\mathbf{0.0381}$ $\mathbf{0.0190}$ $\mathbf{731.25}$
Promotion ($\uparrow$) 28.4% 80.6% 43.4%
KdV Equation. The KdV equation describes the evolution of nonlinear wave phenomena. In this study, due to the absence of one-dimensional instances in the PeRCNN model, we designed our own physical embedding module and corresponding model, following the principles of PeRCNN as detailed in Appendix. We observed significant discrepancies between the predictions of the PeRCNN model and the ground truth, as shown in the blue box of Figure 5(a). Meanwhile, the other baseline models performed poorly at the start, likely due to the complexity of the KdV equation. In contrast, our proposed model demonstrated substantial advantages over the baseline models in terms of predictive accuracy. It maintained a consistent basic shape with the ground truth values, even in long-term forecasts. As shown in Table 1, the PIMRL framework achieved 30% to 50% improvements in the evaluation metrics, highlighting a significant advancement in the field.
2D Burgers Equation. On the right side of Figure 5(b), we observe that besides our framework, only PeRCNN achieved satisfactory results after training on the micro-scale, while the remaining baseline models failed to capture the physical changes effectively. Additionally, in the areas marked by the blue box in Figure 5(b), we can see errors in PeRCNN’s predictions for long-term forecasts, whereas our PIMRL framework demonstrates superior performance, as also visually depicted in Figure 5(b). In the error propagation shown in Figure 5(b), we notice that although initially our framework exhibits higher RMSE compared to the PeRCNN model due to the macro-scale module, the accumulated errors lead to inaccurate predictions by PeRCNN over time. While PeRCNN and PIMRL exhibit the same performance on the HCT metric, this similarity arises from insignificant changes in the later stages. The cumulative error of PeRCNN does not reflect in this particular metric. As shown in Table 1, there are significant improvements in both RMSE and MAE. Additionally, the noticeable divergence of PeRCNN can be observed in the snapshot of the final step.
2D FitzHugh-Nagumo Equation. In the Figure 5(c), we can observe that the purely data-driven method mentioned earlier fails to fully exhibit its original performance on this relatively small dataset. FNO and FNO-coarse represent the effectiveness of the FNO method when trained on datasets of different granularities, respectively. It is worth noting that the performance of FNO-coarse surpasses that of the FNO model trained on micro-scale data. This result effectively validates our hypothesis that error accumulation can significantly impact the model’s performance in long-term multi-step predictions. Quantitatively, our model has improved by at least 10% compared to the currently best-performing model. Furthermore, in the Table 1 we observe that the variation curves of RMSE and other metrics for our model and the PeRCNN model exhibit a striking similarity, yet our model outperforms PeRCNN. This not only indicates that PeRCNN effectively leverages its supervised refinement capabilities but also demonstrates its ability to mitigate the influence of cumulative errors through the evolution of larger time stepping.
2D Gray-Scott Equation. As shown in the left part of Figure 5(d), only PIMRL effectively carried out long-term predictions, showcasing a clear demonstration of the cumulative errors of PeRCNN in this case. The model exhibited significant deviations from ground truth values over extended periods, highlighting its limitations in capturing long-term variations. Among the other models, FNO-Coarse exhibited the best performance. It is evident that FNO performed poorly in the absence of physically informed embeddings, particularly with small time stepping. In this case, Table 1 further validates the superior performance of PIMRL with both RMSE and MAE showing an improvement of over 70%, and the HCT similarity of PIMRL still at 0.99 at the final prediction step. This underscores the effectiveness of the PIMRL framework in enhancing predictive accuracy and maintaining consistency in the predictions throughout the forecasting process. Such robust performance metrics highlight the potential of PIMRL in addressing complex data-driven challenges.
3D Gray-Scott Equation. As shown on the right side of Figure 5(e), the predictions under our framework closely align with the ground truth. Through box plots and error evolution graphs, it is evident that PeRCNN, as the best-performing model among the baselines in our study, outperformed the other models lacking physical knowledge embeddings, especially when compared to FNO trained on the same dataset. In quantitative analysis of Table 1, the improvements are also significant, not only in the overall evaluation metrics of RMSE and MAE in 28.4% and 80.6% but also in the substantial growth of the HCT in 43.4%.
Table 2: Results for ablation study.
Model RMSE
PIMRL w/o Connect 0.1975
FNO-MRL 0.7854
PIMRL w/o Pretraining 0.2599
PIMRL w/o Physics-based FD Conv 0.1738
PeRCNN w/o Physics-based FD Conv NaN
4.2 Ablation Study
To evaluate the impact of components of PIMRL and demonstrate the effectiveness of our framework structure, we have designed five novel models and provided their RMSE results in the FN case. (1) The ablation study with the “PIMRL w/o Connect”, where the connections between the micro-module and the macro-module are removed, is designed to demonstrate the effectiveness of the PIMRL framework’s structural design. This experiment, which leaves only the serial structure of the two modules, shows that the connections within the PIMRL framework are essential for its performance. (2) The “FNO-MRL” replaces the micro-scale module containing physical information, PeRCNN, with the data-driven model FNO, aiming to validate the efficacy of the physical embedding in the micro-scale module within our framework. (3) “PIMRL w/o Pretraining” provides a perspective on the training method by eliminating the pretraining step for the micro module. In this ablation study, the absence of pretraining leads to inferior performance compared to the PIMRL. This demonstrates that directly introducing large time intervals for training can deprive the micro module of the opportunity to learn fine-grained changes, similar to why PeRCNN cannot be directly used with large time intervals. (4) “PIMRL w/o Physics-based FD Conv” indicates the removal of the Physics-based FD Convolution. This ablation study emphasizes the effectiveness of the Physics-based FD Conv by showing the performance degradation when it is omitted. (5) “PeRCNN w/o Physics-based FD Conv” is a version of PeRCNN without the Physics-based Finite Difference (FD) Convolution. While this version can initially make relatively accurate predictions, the errors accumulate over subsequent iterations, eventually becoming unacceptably large.
The five ablation studies conducted not only validate the effectiveness of the interaction design by removing the connection mechanisms but also highlight the contribution of our proposed connection approach. Additionally, by substituting the micro-modules with non-physics-embedded data-driven models, the experiments confirm the efficacy of the physics-embedding within our PIMRL framework. Subsequent pre-training ablation experiments and the removal of the Physics-based FD Conv further substantiate the effectiveness of the corresponding methods and modules.
4.3 Inference Time
Figure 6: Computational time for comparison
In the aforementioned cases, PIMRL not only achieves state-of-the-art performance in long-term predictions but also significantly reduces the computational time cost. Compared to traditional methods such as Direct Numerical Simulation (DNS), our framework is substantially faster, demonstrating a significant improvement in computational efficiency (given the same computing facility, aka, CPU). Additionally, PIMRL not only delivers superior predictive accuracy, but also outperforms PeRCNN in terms of computational efficiency. The numerical methods used for DNS are consistent with the parameter settings during data generation. The difference lies in the required prediction time length, which needs to be the same as that of PIMRL and PeRCNN for a fair comparison.
5 Conclusions
This paper introduces a new multi-scale temporal model named Physics-Informed Multi-Scale Recurrent Learning (PIMRL) for prediction of spatiotemporal dynamics. Adhering to the idea of controlling cumulative error by reducing iterations, PIMRL integrates modules across different scales within its framework to realize this concept, and it can efficiently leverage multi-scale data, which facilitates learning physical laws with limited resources. From fluid dynamics to reaction-diffusion systems, PIMRL demonstrates superior performance in handling multi-scale data, providing accurate long rollout predictions. Our future work will optimize the macro-scale model for time-series tasks to improve computational efficiency and predictive accuracy. We also plan to integrate super-resolution techniques to enhance the model’s adaptability to various spatiotemporal scales, broadening its application to more complex physical systems.
Acknowledgement
The work is supported by the National Natural Science Foundation of China (No. 62276269 and No. 92270118), the Beijing Natural Science Foundation (No. 1232009), and the Fundamental Research Funds for the Central Universities (No. 202230265).
APPENDIX
Appendix A Impact statement
The paper endeavors to devise a adaptable framework that expedites simulations and predictive analyses of physical systems by utilizing multi-scale temporal data. This framework harmoniously integrates data-driven methodologies with physics-informed principles, striking a delicate balance between empirical insights and theoretical underpinnings in its application. The framework can be widely applied in various research fields including material science, turbulent flow prediction, chemical engineering, and so forth. Our research is exclusively conducted for the pursuit of scientific objectives and does not entail any potential ethical concerns or risks.
Appendix B Physical Filter
In the one-dimensional problem KdV, the original paper of PeRCNN did not provide the corresponding model design. Following their concept, we present the corresponding physical filter.
$f_{\text{xxx}}=\frac{-f_{i-2}+2f_{i-1}+0f_{i}-2f_{i+1}+f_{i+2}}{2h^{3}}$ (S1)
Figure S1: Filter for KdV
As shown in the Figure S1, we designed a Physical-Filter to represent $\frac{\partial^{3}u}{\partial x^{3}}$ in the KdV equation. This approach leverages the inherent physical properties of the system to accurately model the third-order spatial derivative, thereby enhancing the accuracy and efficiency of the numerical solution. Among the parameters, $h$ indicates the $\Delta x$ in the cases. The Boundary Padding is an approach to adapt to periodic boundary conditions by replacing the original zero padding with a periodic boundary padding.
Appendix C Implementation Details
C.1 Overview
The overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
Figure S2: Overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
In the main body of this paper, we have elaborated on the architectures at the micro and macro scales. Within the macro-scale module, there are components including an encoder, a decoder, and a Residual Long Short-Term Memory (ResidualLSTM). The following section will provide a detailed exposition of their configuration specifics.
C.2 Encoder and Decoder
In our framework, the autoencoder is employed not for the purpose of minimizing reconstruction loss. Instead, the encoder is utilized to extract features, while the decoder serves to project the output of ConvLSTM into the physical space as residuals. The primary goal of the autoencoder in this context is to map an input to a low-dimensional latent space, and subsequently decode it to the original dimension at the output, facilitating the feature extraction and residual projection processes in the framework.
C.3 ResidualLSTM
The structure of ResidualLSTM has been clearly illustrated in the main text. Here, we will elucidate the architecture of the ConvLSTMcell as the Figure S3.
Figure S3: ConvLSTMCell
$c_{i}$ $=\sigma(\text{W}_{xi}(x_{t})+\text{W}_{hi}(h_{t-1}))$ (S2)
$c_{f}$ $=\sigma(\text{W}_{xf}(x_{t})+\text{W}_{hf}(h_{t-1}))$
$c_{c}$ $=c_{f}\cdot c+c_{i}\cdot\tanh(\text{W}_{xc}(x_{t})+\text{W}_{hc}(h_{t-1}))$
$C_{t}$ $=\sigma(\text{W}_{xo}(x_{t})+\text{W}_{ho}(h_{t-1}))$
$h_{t}$ $=c_{o}\cdot\tanh(c_{c})$
The set of equations S2 presented above outlines the operations of a ConvLSTM cell. The equations involve computations of various gates and states within the cell, including input gate $c_{i}$, forget gate $c_{f}$, cell state $c_{c}$, output $C_{t}$, and the updated hidden state $h_{t}$. These equations govern the flow of information and transformations within the ConvLSTM cell, enabling the model to process spatiotemporal data efficiently by considering spatial dimensions in the calculations. The parameters $\text{W}{xi}$, $\text{W}{hi}$, $\text{W}{xf}$, $\text{W}{hf}$, $\text{W}{xo}$, and $\text{W}{ho}$, depicted in Figure S3, manage information from inputs ($x_{t}$) and history ($h_{t-1}$,$C_{t-1}$) in a convolutional manner.
Appendix D Baseline Models
In order to compare and evaluate the performance of our proposed method, we have trained multiple state-of-the-art (SOTA) baseline models as well as classical models, and compared them with our model. The introductions to each baseline model are presented below, while the training details are outlined in subsequent sections E.
Fourier Neural Operator (FNO) [10]. FNO is a method that combines Fourier transforms with neural networks. This approach comprises two main components. The first component involves performing Fourier transforms on the system state quantities, learning certain information in the frequency domain, and then applying an inverse transform. The second component utilizes convolutions to process the system state quantities, complementing the information not captured during the frequency domain learning. The combination of these two components serves as the final result. We make two FNO models sharing the same architecture, which training in micro and macro scale datasets respectively and inferring in micro and macro time intervals. In the same model, we conducted training separately for two types of data, resulting in FNO and FNO-coarse.
ConvLSTM [35]. ConvLSTM is a specialized neural network architecture that combines convolutional and LSTM layers to effectively model spatial and temporal dependencies in sequential data.
PeRCNN [19]. PeRCNN represents a physics-informed learning methodology, embedding physical laws directly into the neural network architecture. It incorporates multiple parallel convolutional neural networks (CNNs), leveraging the simulation of polynomial equations through feature map multiplication. By doing so, PeRCNN augments the model’s extrapolation and generalization capabilities.
Appendix E Training Details
All experiments were conducted on a single 80GB Nvidia A100 GPU, using an Intel(R) Xeon(R) Platinum 8380 CPU (2.30GHz, 64 cores). We only give some of the changed parameters here, and the other hyperparameters remain the same as the original text.
PIMRL. The micro-module has been trained like PeRCNN in the same way. The architecture of the whole PIMRL model, illustrated in the paper, utilizes the Adam optimizer with a learning rate of $5\times 10^{-3}$. The model have different parameters in different cases.
Additionally, we implement the StepLR scheduler to adjust the learning rate by a factor of 0.98 every 200 epochs. The pretraining details is same to the baseline model PeRCNN.
FNO and FNO-coarse. The network structure of FNO remains largely in line with the original study, with the primary adjustment being the adoption of an autoregressive training method for this model. We employ the Adam optimizer with a learning rate of $1\times 10^{-3}$.
ConvLSTM. For ConvLSTM, we implement the ConvLSTM architecture like the macro-scale module of PIMRL. The StepLR scheduler is utilized with a step size of 200 and a gamma value of 0.98. The optimizer of choice is Adam, featuring a learning rate set at $1\times 10^{-3}$.
PeRCNN. For PeRCNN, the training is different from the original paper since the micro-scale data only get a few pairs of continuous data. It is impossible to train PeRCNN in 400 to 800 steps at the same time. We employ the Adam optimizer with learning rate of $1\times 10^{-3}$ and the StepLR scheduler to adjust the learning rate by a factor of 0.98 per 200 epochs.
Appendix F Evaluation Metrics
In this paper, we have adopted some classical evaluation metrics such as RMSE, MAE and HCT. Root Mean Square Error (RMSE) quantifies the average error magnitude between estimated and actual values, serving as a gauge of the model’s precision. Conversely, Mean Absolute Error (MAE) assesses the average absolute disparity between anticipated and observed values, denoting the true scale of discrepancies.
The definitions of these metrics are as follows:
RMSE (Root Mean Square Error): $\quad\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (S3)
MAE (Mean Absolute Error): $\quad\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
HCT (High Correction Time): $\sum_{i=1}^{N}\Delta t\cdot 1(\text{PCC}(y_{i},\tilde{y}_{i})>0.8)$
In the above equationsS3, $n$ represents the number of trajectories, $y_{i}$ represents the true value, and $\hat{y}_{i}$ represents the predicted value of the model. The PCC is the Pearson correlation coefficient, which is a statistical metric used to measure the linear correlation between two variables.
Appendix G Dataset Informations
The IC for the Kdv equation is created by summing multiple sine waves with random amplitudes, phases, and frequencies, resulting in a complex waveform. Initial conditions for the Burgers’ equation are generated randomly according to a Gaussian distribution. The FN equation is initialized with random Gaussian noise for a warm-up period, after which time sequences are extracted to form the dataset. The GS equation starts the reaction from random initial positions then diffuses.
Table S1: Summary of experimental settings for different cases.(The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training)
Case Numerical Methods Spatial Grid Time Grid Training Trajectories Test Trajectories
Kdv FVM 256 0.01s 5 2
Burgers FDM $128^{2}$ 0.001 13 3
FN FDM $128^{2}$ 0.5 5 3
2DGS FDM $128^{2}$ 0.002 2 3
3DGS FDM $96^{3}$ 0.25 3 2
Korteweg-de Vries Equation. The Korteweg-de Vries system, which elucidates the evolution of waves in nonlinear wave phenomena, can be described by the equation:
$\frac{\partial u}{\partial t}=-u\frac{\partial u}{\partial x}-\frac{\partial^{3}u}{\partial x^{3}}$ (S4)
We got the 8 sets of data: 5 for training, 1 for validation, and 2 for testing. The data sets had spatial domain size $x\in[0,64]$, where $\Delta t$ is 15 times $\delta t$ and $\delta t=0.01s$.
2D Burgers Equation. The 2D Burgers’ equation is commonly employed as a benchmark model for comparing and evaluating different computational algorithms, and describes the complex interaction between nonlinear convection and diffusion processes in the way like:
$\frac{\partial u}{\partial t}$ $=-uu_{x}-vu_{y}+\nu\nabla^{2}u,$ (S5)
$\frac{\partial v}{\partial t}$ $=-uv_{x}-vv_{y}+\nu\nabla^{2}v.$ (S6)
The $u$ and $v$ are the fluid velocities and $\nu$ denotes the viscosity coefficient. In this case, we choose $\nu=0.005$, where $\delta t=0.001s$.
2D FitzHugh-Nagumo Equation. The FitzHugh-Nagumo system can be described by the equation:
$\frac{\partial u}{\partial t}$ $=\mu_{u}\nabla^{2}u+u-u^{3}-v+\alpha,$ (S7)
$\frac{\partial v}{\partial t}$ $=\mu_{v}\nabla^{2}v+(u-v)\beta.$ (S8)
The coefficients $\alpha=0.01$ and $\beta=0.25$, governing the reaction process, take distinct values, while the diffusion coefficients are $\mu_{u}=1$ and $\mu_{v}=100$. In terms of time, $\delta t=0.002s$.
Gray-Scott Equation. The Gray-Scott equations describe the temporal and spatial variations of chemical concentrations in reaction-diffusion systems, which can be described by the equation:
$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+F(1-u),$ (S9)
$\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(F+k)v.$ (S10)
Here, in the two-dimensional case, $D_{u}$ and $D_{v}$ represent the diffusion coefficients of the two substances, with specific values of $D_{u}=2.0\times 10^{-5}$ and $D_{v}=5.0\times 10^{-6}$. $F=0.04$ denotes the growth rate of the substance, while $k=0.06$ signifies its decay rate. In the 2D Gray-Scott case, we got 5 trajectories for training, 1 trajectory for validation and 3 trajectories for testing, where $\delta t=0.5s$.
In the three-dimensional case, we have the parameters: $DA=0.2$, $DB=0.1$, $F=0.025$, and $k=0.055$. We got 3 trajectories for training, 1 trajectory for validation and 2 trajectories for testing, where $\Delta t=15\delta t$ and $\delta t=0.25s$.
Appendix H Supplement results
Considering the constraints on the length of the main text, we have placed some important supplementary results in the appendix of the article. First, we conducted additional experiments on the FN equation, expanding our tests to include more advanced baselines such as F-FNO, U-NO, and MWT (shown in table S2). The experimental results indicate that data-driven methods do not perform well on small training datasets, which further corroborates the effectiveness of the PIMRL proposed in our paper.
To demonstrate the stability of PIMRL, we have included corresponding error bars in the appendix. These results provide sufficient evidence that our model maintains good stability even when trained with a very limited amount of data. The inclusion of error bars offers a quantitative assessment of the variability in our model’s predictions, thereby reinforcing the robustness of PIMRL under data-scarce conditions.
Furthermore, to further validate the capabilities of our model, we have included an additional dataset, introducing the classic 1D Burgers equation dataset from FNO (Fourier Neural Operator). The experimental results show that our model continues to exhibit significant improvements over other baseline models on this dataset. This not only demonstrates the effectiveness of PIMRL in tackling challenging partial differential equation problems but also highlights its broad applicability and superior performance.
Finally, we provide a comparison of the running time, parameter size, and GPU memory usage between PIMRL and the baseline models. This analysis offers insights into the computational efficiency and resource requirements of our model relative to existing approaches, further highlighting the practical advantages of PIMRL in terms of performance optimization and scalability.
Table S2: The Result of U-NO, F-FNO, MWT and FNO in the FN Case.
Metrics U-NO F-FNO MWT FNO
RMSE 0.3675 0.2280 0.3494 0.1878
MAE 0.1465 0.1350 0.2228 0.1634
Table S3: Results with Error Bar under RMSE metric on all Cases.
Model PIMRL PeRCNN
Kdv $0.0457\pm 0.0053$ $0.0942\pm 0.0082$
Burgers $0.0068\pm 0.0006$ $0.0075\pm 0.0008$
2DGS $0.0133\pm 2.4\times 10^{-12}$ $0.0455\pm 1.9\times 10^{-11}$
FN $0.1349\pm 0.0040$ $0.1591\pm 0.0061$
3DGS $0.0381\pm 0.0015$ $0.0532\pm 0.0027$
Table S4: The experiments between PIMRL and baseline models on the Burgers example of FNO benchmark.
Models RMSE MAE
PIMRL 0.0050 0.0039
PeRCNN 0.0137 0.0109
FNO 0.0626 0.0502
FNO-coarse 0.0415 0.0332
Table S5: Running Time, Parameter Size, and GPU Memory for PIMRL, U-NO, MWT, FNO, and ConvLSTM
Model Running Time Parameter Size GPU Memory
PIMRL 9 s 3.33 M 1728 M
U-NO 7 s 15.29 M 2320 M
MWT 12 s 0.09 M 1732 M
FNO 2 s 8.39 M 1580 M
ConvLSTM 5 s 3.32 M 1708 M

### v4

@@
-PIMRL: Physics-Informed Multi-Scale Recurrent Learning for Spatiotemporal Prediction
+PIMRL: Physics-Informed Multi-Scale Recurrent Learning for Burst-Sampled Spatiotemporal Dynamics
 Abstract
-Simulation of spatiotemporal systems governed by partial differential equations is widely applied in fields such as biology, chemistry, aerospace dynamics, and meteorology.
-Traditional numerical methods incur high computational costs due to the requirement of small time steps for accurate predictions.
-While machine learning has reduced these costs, long-term predictions remain challenged by error accumulation, particularly in scenarios with insufficient data or varying time scales, where stability and accuracy are compromised.
-Existing methods often neglect the effective utilization of multi-scale data, leading to suboptimal robustness in predictions.
-To address these issues, we propose a novel multi-scale learning framework, namely, the Physics-Informed Multi-Scale Recurrent Learning (PIMRL), to effectively leverage multi-scale data for spatiotemporal dynamics prediction.
-The PIMRL framework comprises two modules: the micro-scale module embeds physical knowledge into neural networks via pretraining, and the macro-scale module adopts a data-driven approach to learn the temporal evolution of physics in the latent space.
-Experimental results demonstrate that the PIMRL framework consistently achieves state-of-the-art performance across five benchmark datasets ranging from one to three dimensions, showing average improvements of over 9% in both RMSE and MAE evaluation metrics, with maximum enhancements reaching up to 80%.
-1 Introduction
-In the field of natural sciences, physical systems governed by partial differential equations (PDEs) have found widespread applications across disciplines including biology, chemistry, meteorology, etc.
-[1, 2, 3, 4, 5].
-Although numerical methods have been regarded as reliable tools for modeling these systems, the use of Direct Numerical Simulation (DNS) faces significant hurdles due to inherent limitations.
-DNS necessitates high spatial resolution and fine time stepping, resulting in considerable computational demands and prolonged processing times.
-A case in point is the simulation of aerodynamic flows around aircraft, which typically requires the generation of millions of grid points, thereby imposing prohibitive computational requirements [6, 7].
-Moreover, these numerical simulation methods require complete physical prior knowledge, such as PDE formula, parameters, and initial/boundary conditions [8].
-The ongoing development of artificial intelligence (AI) has propelled research into data-driven simulation methods, showcasing significant potential [9, 10, 11, 12].
-These methods, which do not require prior physics knowledge, offer user-friendly solutions.
-They also overcome traditional constraints on resolution and small time stepping, ensuring accurate solutions.
-Nevertheless, when faced with sparse data and multi-scale temporal challenges, these methods often struggle to optimally utilize data.
-They may either discard micro-scale data in favor of macro-scale data or vice versa, leading to compromised accuracy.
-Multi-scale Burst sampling is a technique capturing multiple samples at a high rate over a short period of time to record rapidly changing events or transient phenomena, e.g., fast dynamics, followed by a low sampling rate to capture slow dynamics (see Figure 1).
-Compared to traditional continuous sampling methods, it provides high-resolution data at critical moments while maintaining resource efficiency.
-But conventional uniform-scale models struggle to fully utilize such data.
-The latent ODE approach [13] may work for multi-scale sampled data, but remains unclear whether it can address PDE problems.
-There is an urgent need for a new method to handle such a type of data for spatiotemporal systems.
+Deep learning has shown strong potential in modeling complex spatiotemporal dynamics.
+However, most existing methods depend on densely and uniformly sampled data, which is often unavailable in practice due to sensor and cost limitations.
+In many real-world settings, such as mobile sensing and physical experiments, data are burst-sampled with short high-frequency segments followed by long gaps, making it difficult to learn accurate dynamics from sparse observations.
+To address this issue, we propose Physics-Informed Multi-Scale Recurrent Learning (PIMRL), a novel framework specifically designed for burst-sampled spatiotemporal data.
+PIMRL combines macro-scale latent dynamics inference with micro-scale adaptive refinement guided by incomplete prior information from partial differential equations (PDEs).
+It further introduces a temporal message-passing mechanism to effectively propagate information across burst intervals.
+This multi-scale architecture enables PIMRL to model complex systems accurately even under severe data scarcity.
+We evaluate our approach on five benchmark datasets involving 1D to 3D multi-scale PDEs.
+The results show that PIMRL consistently outperforms state-of-the-art baselines, achieving substantial improvements and reducing errors by up to 80% in the most challenging settings, which demonstrates the clear advantage of our model.
+Our work demonstrates the effectiveness of physics-informed recurrent learning for accurate and efficient modeling of sparse spatiotemporal systems.
+Introduction
 Figure 1: Multi-scale sampling, where $\Delta\tau$ denotes the micro-scale time interval for fast dynamics, $\Delta t$ the macro-scale time interval for slow dynamics, and $\zeta_{t}$ the scale separation variable (typically $\zeta_{t}<1$ or $\zeta_{t}\ll 1$).
-Data-driven methods face significant limitations when addressing challenges related to insufficient data and multiple time scales mentioned above.
-To overcome these obstacles, current approaches integrate physical knowledge into the model learning process.
-Specifically, physics-informed neural networks (PINNs) [14, 15, 16, 17, 18] design initial and boundary conditions as penalty terms in the loss function, thereby leveraging physical laws in a “soft” way.
-However, this soft embedding approach can sometimes lead to unsatisfactory results.
-Also, to some extent, the embedding methods used in PINNs are unable to ensure that the model fully adheres to the embedded physical conditions.
-Directly embedding physical equations into the model architecture, as seen in the physics-encoded recurrent convolutional neural network (PeRCNN) [19], ensures strict adherence to the underlying physical laws.
-This approach addresses the issue of soft constraints in PINNs and enhances the model’s interpretability and generalization capabilities.
-However, this approach requires continuous micro-scale data, which imposes stringent demands on the dataset and can introduce instability in long-term predictions.
-Alternatively, methods such as the learned interpolation (LI) model [20] and multiscale simulation frameworks [21] integrate numerical techniques with neural networks.
-These hybrid approaches enhance simulation efficiency while maintaining acceptable accuracy, although they still have some of the limitations inherent in traditional numerical methods.
-Moreover, we observe that the error increases along with the prediction horizon, a phenomenon that is widespread.
-Although existing efforts are predominantly placed on enhancing the model’s predictive capability, long-term prediction of spatiotemporal dynamics still suffers from error accumulation.
-Hence, properly controlling the accumulation of errors is becomes crucial.
-Based on this core idea, we are motivated to leverage comprehensive multi-scale data through a novel Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework to tackle the problems mentioned above.
-This framework consists of macro-scale and micro-scale modules.
-The micro-scale module, pretrained to learn the underlying physical laws, enhances the accuracy of simulations.
-The macro-scale module reduces the accumulation of errors by minimizing the number of rollout iterations for the micro-scale module, thereby enhancing long-term predictive performance.
-The main contributions of this paper are summarized as follows:
-  • We proposed a new PIMRL framework that effectively leverages information from multi-scale data for long-term spatiotemporal dynamics prediction.
-The concept of reducing the accumulation of errors is achieved through the integration of the macro-scale and micro-scale modules.
-  • We designed a novel message passing mechanism between micro- and macro-scale modules, which effectively transmits physical information, enhances the micro-module’s correction capability, and reduces the number of micro-scale rollout iterations through the macro-scale module.
-  • The PIMRL model achieved optimal performance in effectively predicting the duration of multiple different cases from fluid dynamics to physical systems, demonstrating its scalability and laying a solid foundation for more generalizable models in the future.
-2 Related Work
-Simulation tasks often aim to solve partial differential equations (PDEs) accurately and efficiently.
-Previous researchers have developed numerical methods achieving high precision with many nodes and short time steps.
-To speed up simulations, deep learning techniques have evolved from applying conventional algorithms to designing new models that integrate physical knowledge, including hybrid methods and models embedding physical principles.
-Computational fluid dynamics.
-Computational Fluid Dynamics (CFD) is a branch of fluid mechanics that uses numerical methods and algorithms to analyze and solve problems involving fluid flow [1, 2, 3, 4, 5].
-When faced with complex problems, unacceptable time and computational costs are the primary obstacles limiting the application of CFD.
-Moreover, when the corresponding physical equations contain unknown parameters or even unknown terms, numerical methods are unable to perform accurate simulations.
-The aforementioned methods have inspired us to leverage physical knowledge.
-However, to overcome the aforementioned issues, we have decided to introduce deep learning methods into our PIMRL framework.
+Learning dynamical systems from data is a central challenge in scientific machine learning.
+PDE-governed physical systems are ubiquitous in disciplines such as biology, chemistry, and meteorology (anderson1995computational; darwish2016finite).
+While traditional numerical solvers are reliable, direct numerical simulation (DNS) is often computationally prohibitive.
+It typically requires high spatial and temporal resolutions, long runtimes, and complete knowledge of the governing PDEs, model parameters, and initial/boundary conditions (ferziger2019computational; goc2021large).
+To alleviate the high computational cost of DNS, data-driven deep learning methods such as DeepONet (lu2021learning), FNO (li2020fourier), and DPOT (hao2024dpot; azizzadenesheli2024neural) aim to learn solution operators directly from data.
+These approaches bypass the need for explicit PDEs and enable fast inference by learning mappings between function spaces.
+During rollout, they operate with relatively large time steps, thereby reducing simulation time significantly compared to classical solvers.
+However, their success critically depends on the availability of large, high-quality datasets, which are often expensive to obtain via numerical simulations or physical experiments (kanov2015johns; li2024learning).
+To reduce data dependency, physics-informed learning has emerged as an effective paradigm.
+These methods embed physical knowledge into the training process either softly, by augmenting the loss function with PDE residuals, e.g., PINNs (raissi2019physics), PINO (li2024physics), MCNP (zhang2025monte), or more strictly, by designing model architectures that obey physical rules through discrete approximations, e.g., PeRCNN (Rao_2023; rao2022discovering), P2C2Net (WangEtAl2024NeurIPS), LDSolver (yan2025ldsolver), and TSM (sun2023neural).
+While these approaches improve physical fidelity and reduce labeled data requirements, they still face challenges in long-term forecasting.
+Due to their reliance on local update schemes, they typically require small time steps for numerical stability, which can lead to cumulative errors and increased computational cost over long horizons.
+Since the performance of neural solvers is closely tied to the quality and quantity of training data, the strategy employed during data acquisition plays a critical role.
+However, most existing works assume that data are sampled uniformly or continuously over time.
+This assumption does not hold in many real-world scenarios, where data collection is often constrained by factors such as high experimental costs, limited temporal resolution, and hardware limitations.
+In practical applications including mobile sensing, in situ physical experiments, and computational fluid dynamics simulations, data are frequently gathered in a sparse and non-uniform manner, often across multiple temporal scales.
+Such irregular sampling patterns introduce information gaps and temporal inconsistencies, which can significantly degrade the performance of learning-based models that rely on temporally regular inputs (champion2019discovery).
+Among alternative sampling strategies, burst sampling has recently emerged as a promising approach.
+It captures multiple high-frequency samples within short time intervals to record fast-changing or transient dynamics, while alternating with low-frequency or missing data periods to conserve resources.
+This scheme has demonstrated the ability to capture both rapid and slow dynamics efficiently (see Figure 1).
+However, despite its growing practical relevance, burst-sampled, multi-scale data remains largely unexplored in the context of deep learning for spatiotemporal systems.
+To address this underexplored yet practically important challenge, we propose a framework named Physics-Informed Multi-Scale Recurrent Learning (PIMRL).
+PIMRL is designed to model sparse, non-uniform, and burst-sampled spatiotemporal data by coupling latent-space macro-scale inference with micro-scale physical correction.
+The macro module operates in a learned latent space to perform efficient temporal reasoning and accelerate simulation, while the micro module enforces fine-grained physical consistency using pretrained PDE priors.
+To integrate information across time and resolution scales, we introduce a dedicated cross-scale message passing mechanism that fuses information between the two modules, ensuring stability and accuracy during long-term rollouts.
+Our contributions are summarized as follows:
+  • We propose a Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework that effectively leverages information from multi-scale, burst-sampled data to improve long-term spatiotemporal dynamics prediction.
+  • We develop a cross-scale message passing mechanism that fuses physical information between macro and micro modules, ensuring stable latent state evolution and mitigating error accumulation during long rollouts.
+  • Our framework achieves optimal performance in effectively predicting the dynamics of diverse systems ranging from fluid flows to general physical phenomena, demonstrating scalability and generalizability under data-scarce, irregular sampling regimes.
+Related Work
+Accurate and efficient simulation of partial differential equations (PDEs) is essential for many scientific and engineering tasks, but traditional numerical methods are computationally costly.
+Although deep learning can accelerate simulations, these methods often require extensive physical knowledge or large datasets, resulting in poor performance when such resources are limited.
+In particular, current approaches struggle to model physical systems from sparse measurements without detailed physical priors.
+Direct Numerical Simulation.
+Direct Numerical Simulation (DNS) is a computational method that solves the Navier–Stokes equations directly, capturing all spatial and temporal scales of turbulent flows without turbulence modeling (anderson1995computational; darwish2016finite).
+Although DNS yields highly accurate results, it is extremely computationally expensive.
+This immense computational cost and memory demand often limit DNS to relatively simple or small-scale problems, restricting its practical applicability.
+Furthermore, when the governing equations contain unknown parameters or missing physics, DNS may still struggle to provide reliable predictions.
+To overcome these challenges, PIMRL is specifically designed to efficiently predict outcomes by leveraging limited physical prior knowledge and data.
 Deep Learning Methods.
-With the advancement of AI, the application of AI in physical system simulation tasks has become more diverse and profound.
-For example, classical convolutional neural network models [11, 22] and ResNet [23], U-Net [24] what originally used for image segmentation, graph neural networks [25, 26], and transformer-based models [27, 28, 29, 30] are now being employed.
-Neural operators used for learning mappings between function spaces have also seen significant development in physical simulation tasks, such as DeepONet [9], MWT [31], FNO [10, 32, 33, 34], etc.
-In addition, there are several methods specifically designed for spatiotemporal prediction tasks, such as ConvLSTM [35], PredRNN [36], and TrajGRU [37].
-The traditional deep learning methods mentioned above all require sufficient data for training; insufficient data can lead to poor model performance and severe overfitting issues.
-Moreover, error accumulation is another inevitable problem that limits the application of traditional deep learning methods.
-These limitations restrict the effectiveness and reliability of such models, especially in scenarios requiring long-term predictions or when data is scarce.
-The challenges outlined above have led us to propose a framework that effectively controls error accumulation and adeptly processes multi-scale data.
+With the advancement of AI, its application to physical system simulation has become increasingly diverse and profound.
+For example, classical convolutional neural networks (CNNs) (bar2019learning), ResNet (lu2018beyond), graph neural networks (GNNs) (sanchez2020learning; pfaff2020learning), and transformer-based models (wu2024transolver; hang2024unisolver; janny2023eagle; li2024scalable) have been widely adopted.
+Neural operators for learning mappings between function spaces have also seen rapid development, such as DeepONet (lu2021learning), FNO (li2020fourier; rahman2022u; wen2022u), and PIANO (zhang2024deciphering).
 Physics-Informed Deep Learning Methods.
-To incorporate physical information into models, researchers have devised various methods.
-One category includes physics-inspired methods like PhyCRNet [38], PINN [14, 15, 16, 17, 18], PhyLSTM [39], PhySR [40], and PhyMPGN [41].
-Another category involves physical embedding methods, where explicit physics knowledge is embedded into the model to fully leverage physical principles, like EquNN [42] and PDE-Net [43, 44].
-PeRCNN [19, 45] offers a “hard” encoding mechanism to learn the dynamics of physical systems from limited data.
-This approach leverages prior physics knowledge for predictions, thereby equipping strong predictive capabilities and robust generalization across different initial conditions.
-However, all the aforementioned methods suffer from error accumulation, making it difficult to obtain stable and accurate results in long-term prediction tasks.
-PeRCNN effectively utilizes physical information in a way that can be leveraged within PIMRL.
-Hybrid learning methods.
-In recent years, an emerging research direction has been to integrate deep learning methods with traditional numerical methods.
-By combining classical solvers with deep neural networks, hybrid approaches designed on this principle can achieve acceptable accuracy while operating faster than pure numerical methods.
-Examples include the Learned Interpolation (LI) model [20], numerical discretization learning [46], and P2C2Net [47].
-However, in these hybrid methods, the numerical method component is not involved in the training process, and such approaches also require substantial amounts of data.
-The pioneering work multi-scale simulations of complex systems [21] employs a multi-scale framework for predictions, effectively reducing the required time and computational cost while maintaining good accuracy.
-However, this method still requires ample data for training.
-The idea behind hybrid learning methods is very promising, so we have introduced our own message-passing mechanism within our PIMRL framework.
-3 Methodology
-We propose the PIMRL framework, as illustrated in Figure 2, for spatiotemporal dynamics prediction with a small amount of multi-scale training data
-3.1 Partial Differential Equations simulation task
-Simulation tasks are intimately connected with PDEs, which are essential for describing and simulating physical models, and the time-dependent PDEs are defined as:
-$\mathbf{u}_{t}=\mathcal{F}(t,x,\mathbf{u},\nabla\mathbf{u},\mathbf{u}\cdot\nabla\mathbf{u},\nabla^{2}\mathbf{u},\cdot\cdot\cdot;\mu),$ (1)
-where $\mathbf{u}(x,t)$ denotes the spatiotemporal solution field, $\mathbf{u}_{t}$ the first-order time derivative, $\mathcal{F}(\cdot)$ a linear/nonlinear function, $\nabla$ the Nabla operator, $\nabla^{2}$ the Laplace operator, and $\mu$ the PDE parameter.
-In addition, we define the initial and boundary conditions (ICs, BCs) for this equation, namely: $I[\mathbf{u}](x,t=0)=0$ and $B[\mathbf{u}](x,t)=0$ where $I$ and $B$ indicate IC and BC operators.
-3.2 Overview
-In this study, we aim establish a model capable of predicting the evolution of spatiotemporal dynamics for nonlinear PDE systems at a long-term horizon, given limited training data (e.g., a few trajectories resulted from different initial conditions).
-As outlined in the introduction, we have designed the PIMRL framework to mitigate error accumulation while ensuring that it can effectively capture the intrinsic changes in the physical system, rather than disregarding the physical information over large time intervals.
-Additionally, PIMRL is capable of efficiently utilizing multi-scale data.
-Below, we will provide a detailed exposition of the PIMRL framework’s architecture, the corresponding training methodologies, the macro and micro modules, as well as the boundary padding method that incorporates boundary conditions.
-Figure 2: The overall framework architecture, which integrates physics-informed constraints with deep learning.The initial state of the system is denoted by $\mathbf{u}_{0}$.
+To incorporate physical knowledge, researchers have developed various physics-informed approaches, such as PhyCRNet (ren2022phycrnet), PINN (raissi2019physics), PeSANet (ijcai2025p862), and PDE-Net (long2018pde).
+These methods embed physical laws into deep models to improve prediction accuracy.
+PeRCNN (Rao_2023; rao2022discovering) employs a hard encoding of prior knowledge, enabling strong predictive and generalization capabilities even with limited data.
+However, most existing methods still suffer from error accumulation, which limits their long-term stability and accuracy.
+Notably, PeRCNN’s effective physical encoding strategy can be further leveraged in our proposed PIMRL framework.
+Methodology
+Figure 2: The overall framework architecture, which integrates physics-informed constraints with deep learning.
+The initial state of the system is denoted by $\mathbf{u}_{0}$.
 The state predicted by the micro-module after $k$ iterations, where each iteration occurs at intervals of $\delta t$, is represented by $\mathbf{u}^{micro}_{k\delta t}$.
 The predicted value of the physical state from PIMRL is denoted by $\hat{\mathbf{u}}$.
-3.3 Forecasting Architecture
-As shown in Figure 2, the PIMRL framework comprises two key components: the micro-scale module and the macro-scale module.
-The message-passing mechanism is an interaction between the micro-module and the macro-module.
-The information transmitted includes physical knowledge learned by the micro-module and corrections applied to the macro-module.
-Additionally, the macro-module also passes information to future iterations of the micro-module.
-PIMRL is designed to achieve long-term prediction of spatiotemporal dynamics.
+Partial Differential Equations Simulation Task
+Simulation tasks are closely related to PDEs, which fundamentally describe and model various physical phenomena, with the general form of a time-dependent PDE:
+$\mathbf{u}_{t}=\mathcal{F}(t,x,\mathbf{u},\nabla\mathbf{u},\mathbf{u}\cdot\nabla\mathbf{u},\nabla^{2}\mathbf{u},\ldots;\mu),$ (1)
+where $\mathbf{u}(x,t)$ denotes the spatiotemporal solution field, $\mathbf{u}_{t}$ is the first-order time derivative, $\mathcal{F}(\cdot)$ is a (possibly nonlinear) operator, $\nabla$ represents the gradient (nabla) operator, $\nabla^{2}$ is the Laplacian, and $\mu$ denotes the set of PDE parameters.
+In addition, the initial and boundary conditions (ICs and BCs) are specified as, for example, $I[\mathbf{u}](x,t=0)=0$ and $B[\mathbf{u}](x,t)=0$, where $I[\cdot]$ and $B[\cdot]$ denote the operators for the initial and boundary conditions, respectively.
+Overview
+Figure 3: PIMRL includes two main modules: (a) the micro-module, designed to capture local features and small-scale dynamics; and (b) the macro-module, which captures long-term dependencies and global patterns using residual connections.
+We aim to develop a model capable of predicting the evolution of spatiotemporal dynamics for nonlinear PDE systems over a long-term horizon, based on limited training data (e.g., a few trajectories generated from different initial conditions).
+As outlined in the introduction, we have designed the PIMRL to mitigate error accumulation and ensure that it effectively captures the intrinsic changes in the physical system, rather than overlooking significant physical information over extended time intervals.
+Furthermore, PIMRL is capable of efficiently utilizing multi-scale data.
+As shown in Figure 2, PIMRL consists of two main components: a macro-scale module and a micro-scale module.
+The macro-module performs fast, coarse-grained inference in the latent space with large time steps, accelerating prediction and reducing error accumulation.
+The micro-module conducts fine-grained reasoning and adjusts the latent historical states.
+The message-passing mechanism enables interaction between these modules: the micro-module provides physical knowledge and correction signals to the macro-module, while the macro-module passes updated information to guide future micro-module iterations.
 PIMRL operates in a recursive manner, involving rollouts at the micro-module level, the macro-module level, and the overall framework level.
-Only the output from the macro-module contributes to the final output of the PIMRL framework and is used to compute the loss for training the entire framework, whereas the output from the micro-module serves as intermediate variables within the framework.
-And the message-passing mechanism is described as follows:
-  • Firstly, the micro-module rollout is a simple autoregressive process with time step $\delta t$, where the output at the previous time step serves as the input for the next time step.
+Only the output from the macro-module contributes to the final output of the PIMRL framework and is used to compute the loss for training the entire framework, whereas the output from the micro-module refines and updates the latent historical states.
+The message-passing mechanism we proposed is described as follows:
+  • Firstly, the micro-module rollout is a simple autoregressive process with time step $\delta t$, where the output at the previous time step serves as the input for the next step.
   • When the micro-module is involved in the prediction, for every $k$ steps of micro-module with $\delta t$ like Equation 2, the final output of micro-module is passed to the macro-module, and at this point, the output from the macro-module serves as the output of the entire PIMRL model shown as Equation 3.
@@
   • Finally, there is a PIMRL loop that operates in conjunction with the macro-module rollout.
-After every $N-1$ rollout of the macro-module, the micro-module stops participating in the prediction, and the macro-module performs $N$ steps of autoregressive prediction on its own.
+After every N1N-1 rollout of the macro-module, the micro-module stops participating in the prediction, and the macro-module performs $N$ steps of autoregressive prediction on its own.
 This completes a total of $2N$ rollouts.
 Each output from the macro-module during these $2N$ rollouts serves as the output of PIMRL as depicted in Equation 4.
-We utilize two equations to represent the relationship between the micro-scale module and the macro-scale module, denoted as $F_{micro}$ and $F_{macro}$ respectively, as illustrated in Figure 2.
-The details of this process can be represented using the aforementioned symbols as follows:
-[…]
-[…]
-[…]
-where the variable $\mathbf{u}$ denotes the physical state.
-The relationship between the micro-time step $\delta t$ and the macro-time step $\Delta t$ is given by $\Delta t=k\delta t$, where $k$ is an adjustable parameter determined by the time stepping of different scales in the real data.
-3.4 Training Strategies
-Since the PIMRL model consists of micro- and macro-scale modules, adopting brute-force end-to-end training yields unsatisfactory results (see the ablation study).
-Firstly, we establish a pre-training phase where only the micro-module is trained.
-The purpose of this pre-training phase is to enable the micro-module to effectively learn the dynamics of the physical system and the underlying physical laws, free from the influence of the macro-module.
-In the next phase, referred to as the overall training phase, all modules within the PIMRL framework are engaged in training together.
-During this phase, the micro-module benefits from the parameters pre-trained in the pre-training phase, serving to supervise and correct the macro-module.
-The output from the macro-module serves as the output of the entire PIMRL for computing the loss.
-The details are shown in Appendix.
-Figure 3: PIMRL includes two main modules: (a) the micro-module, designed to capture local features and small-scale dynamics; and (b) the macro-module, which captures long-term dependencies and global patterns using residual connections.
-3.5 Micro-scale Module
-The micro-scale module is designed to learn underlying physical laws that govern the spatiotemporal dynamics from micro-scale data with small time stepping, where we adopt the PeRCNN model [19] with the architecture of $\Pi$-block shown in Figure 3(a).
+Cross-scale Message Passing
+We conceptualize the interaction between the micro-scale module ($F_{\text{micro}}$) and the macro-scale module ($F_{\text{macro}}$) as a cross-scale message passing framework, illustrated in Figure 2.
+The process is governed by the following operations:
+First, during a data-rich burst, the micro-module generates a physically-informed “message” by performing a high-fidelity rollout over $k$ fine steps.
+This message generation process is defined by the Micro-module Loop:
+$\text{Micro-module Loop: }\mathbf{u}^{\text{micro}}_{t+k\delta t}=\underbrace{F_{\text{micro}}(...F_{\text{micro}}}_{\times k}(\mathbf{u}_{t}))$ (2)
+Next, this message, $\mathbf{u}^{\text{micro}}_{t+k\delta t}$, is passed to the macro-module, which consumes it to correct its own state. This state update is described by the Macro-module Loop:
+$\text{Macro-module Loop: }\mathbf{\hat{u}}_{t+2k\delta t}=F_{\text{macro}}(\mathbf{u}^{\text{micro}}_{t+k\delta t})$ (3)
+Finally, after being corrected by the message, the macro-module evolves autoregressively for efficient long-term forecasting.
+This complete, multi-step prediction cycle is summarized by the overall PIMRL process:
+$\text{PIMRL: }\mathbf{\hat{u}}_{t+2N\Delta t}=\underbrace{F_{\text{macro}}(...F_{\text{macro}}}_{\times N}(\mathbf{\hat{u}}_{t+Nk\delta t}))$ (4)
+where the $\mathbf{u}$ denotes the physical state.
+The relationship between the micro-time step $\delta t$ and the macro-time step $\Delta t$ is given by $\Delta t=k\delta t$, where $k$ is an adjustable parameter.
+Micro-scale Module
+The micro-scale module is designed to learn underlying physical laws that govern the spatiotemporal dynamics from micro-scale data with small time stepping, where we adopt the PeRCNN model (Rao_2023) with the architecture of $\Pi$-block shown in Figure 3(a).
 In a forward Euler scheme: $\mathbf{u}_{(k+1)\delta t}=\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})\cdot\delta t+\mathbf{u}_{k\delta t},$ where $\delta t$ denotes that the module predicting in micro-scale time stepping.
 We can then approximate the $\mathcal{F}$ by $\hat{\mathcal{F}}$ described as follows:
-^$\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})=\sum^{N_{c}}_{c=1}W_{c}\cdot\left[\prod_{N_{l}}^{l=1}(K_{c,l}\star\hat{\mathbf{u}}_{k\delta t}+b_{l})\right].$ (5)
+$\hat{\mathcal{F}}(\mathbf{u}_{k\delta t})=\sum^{N_{c}}_{c=1}W_{c}\cdot\left[\prod_{l=1}^{N_{l}}(K_{c,l}\star\hat{\mathbf{u}}_{k\delta t}+b_{l})\right],$ (5)
 where $N_{c}$ denotes the channel count, and $N_{l}$ the total number of parallel convolutional layers.
@@
 For each layer $l$ and channel $c$, $K_{c,l}$ designates the specific filter weight, while $b_{l}$ stands for the bias term of that layer $l$.
-In the context of a $1\times 1$ convolutional layer, $W_{c}$ denotes the weight assigned to the $c^{\text{th}}$ channel, with the bias term being omitted here for the sake of simplicity and brevity.
-When a certain term in the governing PDE remains known (e.g., the diffusion term $\Delta\mathbf{u}$), its discretization can be directly embedded in PeRCNN (called the physics-based Conv layer as shown in Figure 3(a)).
+In the context of a $1\times 1$ convolutional layer, $W_{c}$ denotes the weight assigned to the $c^{\text{th}}$ channel, with the bias term being omitted for the sake of simplicity and brevity.
+When a certain term in the governing PDE remains known (e.g., the Laplace operator $\nabla^{2}\mathbf{u}$), its discretization can be directly embedded in PeRCNN (called the physics-based Conv layer as shown in Figure 3(a)).
 The convolutional kernel in such a layer can be set according to the corresponding finite difference (FD) stencil.
 In essence, the physics-based Conv connection is constructed to incorporate known physical principles, whereas the $\Pi$-block is aimed at capturing the complementary unknown dynamics.
-The details of the physics-based FD Conv are provided in Appendix.
-3.6 Macro-scale Module
-The design of the macro-scale module, as a pivotal component of the PIMRL framework, is meticulously crafted to effectively manage and analyze macro-scale data.
-This type of data often poses unique challenges due to the substantial time spans it encompasses, which in turn lead to significant variations in the underlying physical states captured within the data points.
-These variations might be highly nonlinear and dynamic, making it difficult for traditional, physically motivated modeling methods to accurately capture all the nuances and complexities involved.
-As depicted in Figure 3, our macro-scale module utilizes ConvLSTM block with a residual connection.
-The structure of each block, illustrated in Figure 3(b), consists of a pair of encoders and decoders, along with a ConvLSTM cell shown in Appendix Figure.
-The input to this block undergoes mapping by the encoder, where the feature map serves as the characteristic of the latent space.
-Following this, the ConvLSTM cell simulates the dynamics, and the output is mapped back to the physical space through the decoder.
-Finally, the output, as a residual, is added to the feature from the previous time step to generate the prediction for the next time step.
-More details are provided in Appendix.
-Figure 4: Periodic BC padding.
-3.7 Boundary Condition Padding
-Inspired by PeRCNN [19], we introduce BC hard encoding in both micro-scale and macro-scale modules.
-This encoding method ensures that the feature maps comply with the given BCs during the convolution process while also serving the purpose of padding, which involves filling the feature maps before convolution operations.
-Specifically, in this paper, our case adheres to periodic BCs, and the application of the corresponding padding is illustrated in Figure 4.
-This encoding scheme ingeniously incorporates BCs into the padding, thereby enhancing the accuracy of the prediction.
-4 Experiments
-To validate the effectiveness and versatility of our proposed PIMRL framework, we conducted extensive experiments on a diverse set of fluid dynamics and reaction-diffusion systems equations.
-Specifically, we tested our model on the following cases: the 1D Korteweg-de Vries (KdV) equation, the 2D Burgers equation, and three reaction-diffusion (RD) equations.
-These equations represent a range of physical phenomena with varying degrees of complexity and nonlinearity.
-Our results show that PIMRL consistently outperforms existing methods in terms of accuracy and robustness across these challenging cases.
+The details of the physics-based FD Conv are provided in the Physical Filter part of Appendix A.
+Macro-scale Module
+The macro-scale module serves as the long-term predictive module of PIMRL, specifically engineered to bridge the large temporal gaps inherent in burst-sampled data.
+Instead of operating in the computationally expensive physical space, it performs inference within a compact latent space.
+This design choice is motivated by the need for both computational efficiency and the expressive power to capture complex, long-range dependencies that are intractable for traditional, fine-grained solvers.
+The module is implemented as a residual ConvLSTM autoencoder (Figure 3; Appendix B), with components operating together within a message-passing framework.
+An encoder maps the input to a latent space, where the ConvLSTM cell propagates the system’s state forward in time with large, efficient steps ($\Delta t$).
+Crucially, this latent-space evolution is not a pure black-box rollout.
+It is periodically corrected by physics-informed messages from the micro-module.
+This mechanism allows the macro-module to anchor its latent trajectory to physical reality, preventing the error accumulation and drift that plague purely data-driven recurrent models during long-term forecasting.
+By integrating the efficiency of latent-space dynamics with the physical fidelity of micro-scale corrections, our macro-scale module achieves a robust and accurate forecasting capability.
+Case Numerical Methods Spatial Grid Macro Step Micro Step Training Trajectories Test Trajectories
+KdV FVM 256 0.01s 0.01s 5 2
+Burgers FDM $128^{2}$ 0.001s 0.01s 13 3
+FN FDM $128^{2}$ 0.5s 0.01s 5 3
+2D GS FDM $128^{2}$ 0.002s 0.01s 2 3
+3D GS FDM $96^{3}$ 0.25s 0.01s 3 2
+Table 1: Summary of experimental settings for different cases. The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training.
+Experiments
+To validate the effectiveness of our proposed PIMRL framework, we conducted extensive experiments on a diverse set of fluid dynamics and reaction-diffusion systems.
+Figure 4: Error propagation curves and final prediction plots for the KdV case, comparing the PIMRL and baseline models.
+Figure 5: Figures (a)–(c) summarize qualitative comparisons with baselines on Burgers, FN, and 2D GS, showing error propagation and final predictions.
 Datasets.
-We conducted experiments on five datasets, including: Korteweg-de Vries (KdV), 2D Burgers, FitzHugh-Nagumo (FN), 2D Gray-Scott (2D GS), and 3D Gray-Scott (3D GS).
+We conducted experiments on five datasets, including Korteweg-de Vries (KdV), Burgers, FitzHugh-Nagumo (FN), 2D Gray-Scott (2D GS), and 3D Gray-Scott (3D GS).
 The IC for the KdV equation is created by summing multiple sine waves with random amplitudes, phases, and frequencies, resulting in a complex waveform.
@@
 Additionally, we have two sets of data with different time scales originating from the same ICs.
-The micro-scale data $\textbf{U}^{\text{micro}}=\{\textbf{u}_{0},\textbf{u}_{\delta t},\textbf{u}_{2\delta t}\cdot\cdot\cdot\cdot\}\in\text{R}^{\text{micro}}$ is characterized by short and scattered continuous time intervals, while the macro-scale data $\textbf{U}^{\text{macro}}=\{\textbf{u}_{0},\textbf{u}_{\Delta t},\textbf{u}_{2\Delta t}\cdot\cdot\cdot\cdot\textbf{u}_{\text{Tend}}\}\in\text{R}^{\text{macro}}$ exhibits persistent continuity until the end.
+The micro-scale data $\mathbf{U}^{\text{micro}}=\{\mathbf{u}_{0},\mathbf{u}_{\delta t},\mathbf{u}_{2\delta t}\cdot\cdot\cdot\}\in\mathbf{R}^{\text{micro}}$ is characterized by short and scattered continuous time intervals, while the macro-scale data $\mathbf{U}^{\text{macro}}=\{\mathbf{u}_{0},\mathbf{u}_{\Delta t},\mathbf{u}_{2\Delta t}\cdot\cdot\cdot\mathbf{u}_{T_{\text{end}}}\}\in\mathbf{R}^{\text{macro}}$ exhibits persistent continuity until the end.
 The validation and test sets are established based on different ICs but with the same parameters, making it more challenging than extrapolation under the same ICs.
-The dataset configuration is presented in Appendix.
-Figure 5: An overview of the comparison between our PIMRL framework and baseline models is provided, including error propagation curves (left), error boxplots (middle), and final prediction plots (right).
-Figures (a) through (e) respectively showcase the qualitative results for the KdV, Burgers, 2D GS, FN, and 3D GS cases.
-Model training.
-The primary objective is to first pretrain the micro-scale module using the micro-scale data to learn the underlying physics.
-Subsequently, the pretrained micro-scale module is integrated into the overall framework, and the entire model is trained using macro-scale data to capture the spatial evolution patterns over long time stepping.
-Both the pretraining and training processes can be formulated as auto-regressive rollout problems.
-The loss function is defined as $L(\boldsymbol{\theta})=\frac{1}{BN}\sum_{i=1}^{B}\sum_{j=1}^{N}(y_{i,j}-\hat{y}_{i,j})^{2}$, where the $B$ and $N$ denote the number of batches and the batch size.
-The $\boldsymbol{\theta}$ indicates all the trainable parameters and $(y_{i,j}-\hat{y}_{i,j})$ means the difference between the rollout-prediction $\hat{y}_{i,j}$ of $j$-th sample in the $i$-th batch and the corresponding labeled data $y_{i,j}$.
-More training details are shown in Appendix.
 Baseline Models.
 To validate the effectiveness of the proposed PIMRL framework, we introduced several baseline models.
-Firstly, we considered the widely recognized high-performing data-driven model FNO [10], which has been trained on datasets with two different time intervals, denoted as FNO (trained on fine-scale data with small time steps) and FNO-coarse (trained on coarse-scale data with large time steps).
-In particular, we examined the impact of cumulative error on long-term predictions.
-Our results show that the influence of error accumulation is significant.
-Detailed analysis is provided in Section 4.1.
-Secondly, we included the PeRCNN model [19], which embeds physical knowledge in a hard way and demonstrates excellent performance on multiple datasets.
-Due to constraints on the time stepping of the model, PeRCNN is trained only on datasets with small time stepping.
-Lastly, we incorporated ConvLSTM [35], a classic sequential prediction model trained on datasets with large time stepping, serving as the macro-scale module within our framework in a residual way.
-The details are shown in the Appendix.
+Firstly, we considered the widely recognized high-performing data-driven model FNO (li2020fourier), which has been trained on datasets with two different time intervals, denoted as FNO (trained on fine-scale data with small time steps) and FNO-coarse (trained on coarse-scale data with large time steps).
+Secondly, we included the PeRCNN model (Rao_2023), which embeds physical knowledge in a hard way.
+Due to constraints on the time stepping of the model, PeRCNN is trained on datasets with small time steps.
+Lastly, we incorporated CROP (gao2025discretizationinvariance)(CROP-coarse trained on coarse-scale data with large time steps), a method proposed to address problems with generalization and discretization mismatch errors in existing neural operators across different data resolutions.
+Since the original CROP paper did not include 1D and 3D cases, we employ AE-ConvLSTM (vlachas2022multiscale) as a substitute.
+The details are shown in Appendix C.
 Evaluation Metrics.
-To comprehensively evaluate the performance of our model, we adopted several metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and High Correction Time (HCT).
-These metrics provide a multi-faceted assessment of the model’s accuracy and reliability.
+To comprehensively evaluate the performance of our model, we adopted metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and High Correction Time (HCT).
 RMSE is calculated on the macro-scale data to facilitate comparisons between models operating at different granularities.
-MAE provides a measure of the average absolute difference between the predicted and actual values,less sensitive to outliers compared to RMSE.
-HCT evaluates the time it takes for the model to correct its predictions to a high level of accuracy, which is particularly important for long-term prediction.
-Detailed formulas for these metrics are provided in Appendix.
-4.1 Main Results
-Figure 5 illustrates the performance results of our framework in comparison with various baseline models across different cases, including error propagation, showcasing box plots of different quantities and their prediction outcomes.
-Additionally, Table 1 presents relevant quantitative metrics as an expression of the results, where * in Table 1 indicates that the inference process has reached the end of the test data.
-Table 1: Quantitative results of our model and baselines.
+MAE provides a measure of the average absolute difference between the predicted and actual values.
+HCT evaluates the time it takes for the model to correct its predictions to a high level of accuracy.
+Detailed formulas for these metrics are provided in the Evaluation Metrics part of Appendix E.
 Case Model RMSE $\downarrow$ MAE $\downarrow$ HCT (s) $\uparrow$
-KdV FNO 0.4891 0.3300 0.45
+KdV ConvLSTM 5.8507 7.6036 9.6
+FNO 0.4891 0.3300 0.45
 FNO-coarse 0.5461 0.4167 7.8
@@
 PIMRL(Ours) $\mathbf{0.0457}$ $\mathbf{0.0607}$ $\mathbf{46.2}$
-Promotion ($\uparrow$) 51.5% 35.5% 54.0%
-Burgers FNO 0.1561 0.1301 0.104
+Promotion 51.5% 35.5% 54.0%
+Burgers CROP-coarse 0.1103 0.0887 0.064
+FNO 0.1561 0.1301 0.104
 FNO-coarse 0.1094 0.0879 0.064
@@
 PIMRL(Ours) $\mathbf{0.0068}$ $\mathbf{0.0049}$ $\mathbf{3.216}$
-Promotion ($\uparrow$) 9.3% 15.6% 0%
-FN FNO 937 2393980 1.65
+Promotion 9.3% 15.6% 0%
+FN CROP-coarse 0.3950 0.2734 3.03
+FNO 937 2393980 1.65
 FNO-coarse 0.1878 0.1643 4.98
@@
 PIMRL(Ours) $\mathbf{0.1349}$ $\mathbf{0.0990}$ $\mathbf{7.74}$
-Promotion ($\uparrow$) 15.2% 13.1% 10.7%
-2DGS FNO NaN NaN 810
+Promotion 15.2% 13.1% 10.7%
+2D GS CROP-coarse 0.1027 0.0579 1260
+FNO NaN NaN 810
 FNO-coarse 0.0884 0.0629 1335
@@
 PIMRL(Ours) $\mathbf{0.0133}$ $\mathbf{0.0072}$ $\mathbf{1965^{*}}$
-Promotion ($\uparrow$) 70.8% 73.1% 42.5%
-3DGS FNO 0.2798 0.1950 112.5
+Promotion 70.8% 73.1% 42.5%
+3D GS ConvLSTM 0.2081 0.2009 56.5
+FNO 0.2798 0.1950 112.5
 FNO-coarse 0.1042 0.0611 360
@@
 PIMRL(Ours) $\mathbf{0.0381}$ $\mathbf{0.0190}$ $\mathbf{731.25}$
-Promotion ($\uparrow$) 28.4% 80.6% 43.4%
-KdV Equation.
-The KdV equation describes the evolution of nonlinear wave phenomena.
-In this study, due to the absence of one-dimensional instances in the PeRCNN model, we designed our own physical embedding module and corresponding model, following the principles of PeRCNN as detailed in Appendix.
-We observed significant discrepancies between the predictions of the PeRCNN model and the ground truth, as shown in the blue box of Figure 5(a).
-Meanwhile, the other baseline models performed poorly at the start, likely due to the complexity of the KdV equation.
-In contrast, our proposed model demonstrated substantial advantages over the baseline models in terms of predictive accuracy.
-It maintained a consistent basic shape with the ground truth values, even in long-term forecasts.
-As shown in Table 1, the PIMRL framework achieved 30% to 50% improvements in the evaluation metrics, highlighting a significant advancement in the field.
+Promotion 28.4% 80.6% 43.4%
+Table 2: Quantitative results of our model and baselines, where * denotes that the prediction lasts to the end.
+Main Results
+Figure 5 compares our framework with baseline models across cases using box plots of predicted quantities. Table 2 reports the corresponding quantitative metrics.
+1D KdV Equation.
+The KdV equation governs nonlinear wave evolution.
+Notably, significant discrepancies between PeRCNN predictions and ground truth emerge in the blue box of Figure 4, whereas other baselines exhibit poor initial performance attributable to KdV complexity.
+Conversely, PIMRL demonstrates substantial advantages in predictive accuracy over baseline models, achieving 30%–50% improvements in evaluation metrics (Table 2) and thereby highlighting significant field advancement.
 2D Burgers Equation.
-On the right side of Figure 5(b), we observe that besides our framework, only PeRCNN achieved satisfactory results after training on the micro-scale, while the remaining baseline models failed to capture the physical changes effectively.
-Additionally, in the areas marked by the blue box in Figure 5(b), we can see errors in PeRCNN’s predictions for long-term forecasts, whereas our PIMRL framework demonstrates superior performance, as also visually depicted in Figure 5(b).
-In the error propagation shown in Figure 5(b), we notice that although initially our framework exhibits higher RMSE compared to the PeRCNN model due to the macro-scale module, the accumulated errors lead to inaccurate predictions by PeRCNN over time.
-While PeRCNN and PIMRL exhibit the same performance on the HCT metric, this similarity arises from insignificant changes in the later stages.
-The cumulative error of PeRCNN does not reflect in this particular metric.
-As shown in Table 1, there are significant improvements in both RMSE and MAE.
-Additionally, the noticeable divergence of PeRCNN can be observed in the snapshot of the final step.
+Figure 5(a) demonstrates that only PeRCNN and our framework attain satisfactory post-training results on micro-scale data, while other baselines fail to capture physical changes.
+In the blue-boxed regions, PeRCNN shows significant long-term forecast errors compared to PIMRL’s sustained accuracy.
+Corresponding quantitative improvements in RMSE and MAE are documented in Table 2, with PeRCNN’s divergence further confirmed at the final time step.
 2D FitzHugh-Nagumo Equation.
-In the Figure 5(c), we can observe that the purely data-driven method mentioned earlier fails to fully exhibit its original performance on this relatively small dataset.
-FNO and FNO-coarse represent the effectiveness of the FNO method when trained on datasets of different granularities, respectively.
-It is worth noting that the performance of FNO-coarse surpasses that of the FNO model trained on micro-scale data.
-This result effectively validates our hypothesis that error accumulation can significantly impact the model’s performance in long-term multi-step predictions.
-Quantitatively, our model has improved by at least 10% compared to the currently best-performing model.
-Furthermore, in the Table 1 we observe that the variation curves of RMSE and other metrics for our model and the PeRCNN model exhibit a striking similarity, yet our model outperforms PeRCNN.
-This not only indicates that PeRCNN effectively leverages its supervised refinement capabilities but also demonstrates its ability to mitigate the influence of cumulative errors through the evolution of larger time stepping.
+In Figure 5(b), the purely data-driven method fails to achieve its original performance on this relatively small dataset.
+Quantitatively, our model outperforms the best existing model by at least 10%.
+Additionally, Table 2 shows that while the RMSE and other metric curves of our model and PeRCNN are similar, our model consistently achieves better results.
 2D Gray-Scott Equation.
-As shown in the left part of Figure 5(d), only PIMRL effectively carried out long-term predictions, showcasing a clear demonstration of the cumulative errors of PeRCNN in this case.
-The model exhibited significant deviations from ground truth values over extended periods, highlighting its limitations in capturing long-term variations.
-Among the other models, FNO-Coarse exhibited the best performance.
-It is evident that FNO performed poorly in the absence of physically informed embeddings, particularly with small time stepping.
-In this case, Table 1 further validates the superior performance of PIMRL with both RMSE and MAE showing an improvement of over 70%, and the HCT similarity of PIMRL still at 0.99 at the final prediction step.
-This underscores the effectiveness of the PIMRL framework in enhancing predictive accuracy and maintaining consistency in the predictions throughout the forecasting process.
-Such robust performance metrics highlight the potential of PIMRL in addressing complex data-driven challenges.
+As shown in the left part of Figure 5(c), only PIMRL effectively carried out long-term predictions, showcasing a clear demonstration of the cumulative errors of PeRCNN in this case.
+In this case, Table 2 further validates the superior performance of PIMRL with both RMSE and MAE showing an improvement of over 70%, and the HCT similarity of PIMRL at 0.99 at the final prediction step.
+This highlights the effectiveness of PIMRL in improving predictive accuracy and consistency throughout forecasting, demonstrating its potential for tackling complex challenges.
+Figure 6: Error propagation curves and final prediction plots for the 3D GS, comparing the PIMRL and baseline models.
 3D Gray-Scott Equation.
-As shown on the right side of Figure 5(e), the predictions under our framework closely align with the ground truth.
+As shown on the right side of Figure 6, the predictions under our framework closely align with the ground truth.
 Through box plots and error evolution graphs, it is evident that PeRCNN, as the best-performing model among the baselines in our study, outperformed the other models lacking physical knowledge embeddings, especially when compared to FNO trained on the same dataset.
 In quantitative analysis of Table 2, the improvements are also significant, not only in the overall evaluation metrics of RMSE and MAE in 28.4% and 80.6% but also in the substantial growth of the HCT in 43.4%.
-Table 2: Results for ablation study.
+Ablation Study
+We designed five novel models and report them in Table 3.
 Model RMSE
@@
 PIMRL w/o Physics-based FD Conv 0.1738
-PeRCNN w/o Physics-based FD Conv NaN
-4.2 Ablation Study
-To evaluate the impact of components of PIMRL and demonstrate the effectiveness of our framework structure, we have designed five novel models and provided their RMSE results in the FN case.
+PIMRL 0.1349
+Table 3: Results for ablation study in FN-case.
 (1) The ablation study with the “PIMRL w/o Connect”, where the connections between the micro-module and the macro-module are removed, is designed to demonstrate the effectiveness of the PIMRL framework’s structural design.
-This experiment, which leaves only the serial structure of the two modules, shows that the connections within the PIMRL framework are essential for its performance.
+This experiment, which leaves only the serial structure of the two modules, shows that the connections within the PIMRL framework are essential.
 (2) The “FNO-MRL” replaces the micro-scale module containing physical information, PeRCNN, with the data-driven model FNO, aiming to validate the efficacy of the physical embedding in the micro-scale module within our framework.
@@
 (5) “PeRCNN w/o Physics-based FD Conv” is a version of PeRCNN without the Physics-based Finite Difference (FD) Convolution.
-While this version can initially make relatively accurate predictions, the errors accumulate over subsequent iterations, eventually becoming unacceptably large.
-The five ablation studies conducted not only validate the effectiveness of the interaction design by removing the connection mechanisms but also highlight the contribution of our proposed connection approach.
-Additionally, by substituting the micro-modules with non-physics-embedded data-driven models, the experiments confirm the efficacy of the physics-embedding within our PIMRL framework.
-Subsequent pre-training ablation experiments and the removal of the Physics-based FD Conv further substantiate the effectiveness of the corresponding methods and modules.
-4.3 Inference Time
-Figure 6: Computational time for comparison
-In the aforementioned cases, PIMRL not only achieves state-of-the-art performance in long-term predictions but also significantly reduces the computational time cost.
-Compared to traditional methods such as Direct Numerical Simulation (DNS), our framework is substantially faster, demonstrating a significant improvement in computational efficiency (given the same computing facility, aka, CPU).
-Additionally, PIMRL not only delivers superior predictive accuracy, but also outperforms PeRCNN in terms of computational efficiency.
-The numerical methods used for DNS are consistent with the parameter settings during data generation.
-The difference lies in the required prediction time length, which needs to be the same as that of PIMRL and PeRCNN for a fair comparison.
-5 Conclusions
-This paper introduces a new multi-scale temporal model named Physics-Informed Multi-Scale Recurrent Learning (PIMRL) for prediction of spatiotemporal dynamics.
-Adhering to the idea of controlling cumulative error by reducing iterations, PIMRL integrates modules across different scales within its framework to realize this concept, and it can efficiently leverage multi-scale data, which facilitates learning physical laws with limited resources.
-From fluid dynamics to reaction-diffusion systems, PIMRL demonstrates superior performance in handling multi-scale data, providing accurate long rollout predictions.
-Our future work will optimize the macro-scale model for time-series tasks to improve computational efficiency and predictive accuracy.
-We also plan to integrate super-resolution techniques to enhance the model’s adaptability to various spatiotemporal scales, broadening its application to more complex physical systems.
-Acknowledgement
-The work is supported by the National Natural Science Foundation of China (No.
+Addition Results
+To comprehensively evaluate PIMRL, we conducted a series of supplementary experiments, the details of which are provided in Appendix G.
+In particular, we assessed computational efficiency by comparing runtime, parameter count, and memory usage with other physics-embedded methods, and examined the model’s scalability and practicality in various scenarios.
+Furthermore, we investigated the robustness of PIMRL under different training set sizes and analyzed the reliability of its predictions using error bar analyses.
+These additional experiments collectively demonstrate the promising efficiency, scalability, and robustness of PIMRL.
+Conclusion
+In this paper, we introduced the Physics-Informed Multi-Scale Recurrent Learning (PIMRL) framework, the first, to our knowledge, specifically designed to predict the evolution of spatiotemporal systems from multi-scale, burst-sampled data.
+Its novel architecture synergistically combines a macro-scale module for efficient long-term forecasting with a physics-informed micro-scale module for periodic correction, a design that effectively mitigates cumulative error while capturing underlying physical laws from extremely limited observations.
+Rigorous evaluations confirmed its substantial performance gains over state-of-the-art baselines, including a reduction in RMSE of up to 70% on the 2D Gray-Scott dataset, thereby validating the effectiveness of this multi-scale, physics-informed approach.
+While we acknowledge the framework’s current limitations, such as the yet unexplored sensitivity to the duration of data gaps and the presence of input noise, as detailed in Appendix H, we view them as clear and important avenues for future research.
+Going forward, we will improve the macro-module’s efficiency and integrate advanced super-resolution to better exploit the model’s potential, enhance generalization, and widen its impact in scientific computing.
+Acknowledgments
+The work is supported by the Beijing Natural Science Foundation (No.
+1232009) and the National Natural Science Foundation of China (No.
 62276269 and No.
-92270118), the Beijing Natural Science Foundation (No.
-1232009), and the Fundamental Research Funds for the Central Universities (No.
-202230265).
-APPENDIX
-Appendix A Impact statement
+62506367).
+R.Z.
+would like to acknowledge the supported by the China Postdoctoral Science Foundation under Grant Number 2025M771582 and the Postdoctoral Fellowship Program of CPSF under Grant Number GZB20250408.
+Additional experimental details and full appendices are available in the arXiv version of this paper at https://arxiv.org/abs/2503.10253.
+Impact statement
 The paper endeavors to devise a adaptable framework that expedites simulations and predictive analyses of physical systems by utilizing multi-scale temporal data.
@@
 Our research is exclusively conducted for the pursuit of scientific objectives and does not entail any potential ethical concerns or risks.
-Appendix B Physical Filter
+A: Physical Filter
 In the one-dimensional problem KdV, the original paper of PeRCNN did not provide the corresponding model design. Following their concept, we present the corresponding physical filter.
-$f_{\text{xxx}}=\frac{-f_{i-2}+2f_{i-1}+0f_{i}-2f_{i+1}+f_{i+2}}{2h^{3}}$ (S1)
-Figure S1: Filter for KdV
-As shown in the Figure S1, we designed a Physical-Filter to represent $\frac{\partial^{3}u}{\partial x^{3}}$ in the KdV equation.
+$f_{\text{xxx}}=\frac{-f_{i-2}+2f_{i-1}+0f_{i}-2f_{i+1}+f_{i+2}}{2h^{3}}$ (6)
+As shown in the Figure LABEL:filter4kdv, we designed a Physical-Filter to represent $\frac{\partial^{3}u}{\partial x^{3}}$ in the KdV equation.
 This approach leverages the inherent physical properties of the system to accurately model the third-order spatial derivative, thereby enhancing the accuracy and efficiency of the numerical solution.
@@
 The Boundary Padding is an approach to adapt to periodic boundary conditions by replacing the original zero padding with a periodic boundary padding.
-Appendix C Implementation Details
-C.1 Overview
+B: Implementation Details
+Overview
 The overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
-Figure S2: Overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
+Figure 7: Overview of the PIMRL framework, which includes a pretraining stage using micro-scale data for physics-informed Learning and the utilization of a micro-module, informed by learned physics knowledge, to correct the macro-module during training on macro-scale data.
 In the main body of this paper, we have elaborated on the architectures at the micro and macro scales.
@@
 The following section will provide a detailed exposition of their configuration specifics.
-C.2 Encoder and Decoder
+Encoder and Decoder
 In our framework, the autoencoder is employed not for the purpose of minimizing reconstruction loss.
@@
 The primary goal of the autoencoder in this context is to map an input to a low-dimensional latent space, and subsequently decode it to the original dimension at the output, facilitating the feature extraction and residual projection processes in the framework.
-C.3 ResidualLSTM
-The structure of ResidualLSTM has been clearly illustrated in the main text. Here, we will elucidate the architecture of the ConvLSTMcell as the Figure S3.
-Figure S3: ConvLSTMCell
-$c_{i}$ $=\sigma(\text{W}_{xi}(x_{t})+\text{W}_{hi}(h_{t-1}))$ (S2)
+ResidualLSTM
+The structure of ResidualLSTM has been clearly illustrated in the main text. Here, we will elucidate the architecture of the ConvLSTMcell as the Figure 8.
+Figure 8: ConvLSTMCell
+$c_{i}$ $=\sigma(\text{W}_{xi}(x_{t})+\text{W}_{hi}(h_{t-1}))$ (7)
 $c_{f}$ $=\sigma(\text{W}_{xf}(x_{t})+\text{W}_{hf}(h_{t-1}))$
@@
 $h_{t}$ $=c_{o}\cdot\tanh(c_{c})$
-The set of equations S2 presented above outlines the operations of a ConvLSTM cell.
+The set of equations 7 presented above outlines the operations of a ConvLSTM cell.
 The equations involve computations of various gates and states within the cell, including input gate $c_{i}$, forget gate $c_{f}$, cell state $c_{c}$, output $C_{t}$, and the updated hidden state $h_{t}$.
 These equations govern the flow of information and transformations within the ConvLSTM cell, enabling the model to process spatiotemporal data efficiently by considering spatial dimensions in the calculations.
-The parameters $\text{W}{xi}$, $\text{W}{hi}$, $\text{W}{xf}$, $\text{W}{hf}$, $\text{W}{xo}$, and $\text{W}{ho}$, depicted in Figure S3, manage information from inputs ($x_{t}$) and history ($h_{t-1}$,$C_{t-1}$) in a convolutional manner.
-Appendix D Baseline Models
+The parameters $\text{W}{xi}$, $\text{W}{hi}$, $\text{W}{xf}$, $\text{W}{hf}$, $\text{W}{xo}$, and $\text{W}{ho}$, depicted in Figure 8, manage information from inputs ($x_{t}$) and history ($h_{t-1}$$C_{t-1}$) in a convolutional manner.
+Metrics U-NO F-FNO MWT FNO
+RMSE 0.3675 0.2280 0.3494 0.1878
+MAE 0.1465 0.1350 0.2228 0.1634
+Table 4: The Result of U-NO, F-FNO, MWT and FNO in the FN Case.
+C: Baseline Models
 In order to compare and evaluate the performance of our proposed method, we have trained multiple state-of-the-art (SOTA) baseline models as well as classical models, and compared them with our model.
-The introductions to each baseline model are presented below, while the training details are outlined in subsequent sections E.
-Fourier Neural Operator (FNO) [10].
+The introductions to each baseline model are presented below, while the training details are outlined in subsequent sections D: Training Details.
+Fourier Neural Operator (FNO) (li2020fourier).
 FNO is a method that combines Fourier transforms with neural networks.
@@
 In the same model, we conducted training separately for two types of data, resulting in FNO and FNO-coarse.
-ConvLSTM [35]. ConvLSTM is a specialized neural network architecture that combines convolutional and LSTM layers to effectively model spatial and temporal dependencies in sequential data.
-PeRCNN [19].
+ConvLSTM (vlachas2022multiscale).
+ConvLSTM is a specialized neural network architecture that combines convolutional and LSTM layers to effectively model spatial and temporal dependencies in sequential data.
+We use AE-ConvLSTM as our base model.
+PeRCNN (Rao_2023).
 PeRCNN represents a physics-informed learning methodology, embedding physical laws directly into the neural network architecture.
@@
 By doing so, PeRCNN augments the model’s extrapolation and generalization capabilities.
-Appendix E Training Details
+D: Training Details
 All experiments were conducted on a single 80GB Nvidia A100 GPU, using an Intel(R) Xeon(R) Platinum 8380 CPU (2.30GHz, 64 cores).
@@
 We employ the Adam optimizer with learning rate of $1\times 10^{-3}$ and the StepLR scheduler to adjust the learning rate by a factor of 0.98 per 200 epochs.
-Appendix F Evaluation Metrics
-In this paper, we have adopted some classical evaluation metrics such as RMSE, MAE and HCT.
+E: Evaluation Metrics
+We have adopted some classical evaluation metrics such as RMSE, MAE and HCT.
 Root Mean Square Error (RMSE) quantifies the average error magnitude between estimated and actual values, serving as a gauge of the model’s precision.
@@
 The definitions of these metrics are as follows:
-RMSE (Root Mean Square Error): $\quad\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (S3)
+RMSE (Root Mean Square Error): $\quad\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (8)
 MAE (Mean Absolute Error): $\quad\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$
-HCT (High Correction Time): $\sum_{i=1}^{N}\Delta t\cdot 1(\text{PCC}(y_{i},\tilde{y}_{i})>0.8)$
-In the above equationsS3, $n$ represents the number of trajectories, $y_{i}$ represents the true value, and $\hat{y}_{i}$ represents the predicted value of the model.
+[…]
+In the above equations8, $n$ represents the number of trajectories, $y_{i}$ represents the true value, and $\hat{y}_{i}$ represents the predicted value of the model.
 The PCC is the Pearson correlation coefficient, which is a statistical metric used to measure the linear correlation between two variables.
-Appendix G Dataset Informations
+F: Dataset Informations
 The IC for the Kdv equation is created by summing multiple sine waves with random amplitudes, phases, and frequencies, resulting in a complex waveform.
@@
 The GS equation starts the reaction from random initial positions then diffuses.
-Table S1: Summary of experimental settings for different cases.(The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training)
+Model Running Time Parameter Size GPU Memory
+PIMRL 9 s 3.33 M 1728 M
+U-NO 7 s 15.29 M 2320 M
+MWT 12 s 0.09 M 1732 M
+FNO 2 s 8.39 M 1580 M
+ConvLSTM 5 s 3.32 M 1708 M
+Table 5: Running Time, Parameter Size, and GPU Memory for PIMRL, U-NO, MWT, FNO, and ConvLSTM
 Case Numerical Methods Spatial Grid Time Grid Training Trajectories Test Trajectories
@@
 3DGS FDM $96^{3}$ 0.25 3 2
+Table 6: Summary of experimental settings for different cases.(The 3D GS case is downsampled from $96^{3}$ to $48^{3}$ during training)
 Korteweg-de Vries Equation. The Korteweg-de Vries system, which elucidates the evolution of waves in nonlinear wave phenomena, can be described by the equation:
-$\frac{\partial u}{\partial t}=-u\frac{\partial u}{\partial x}-\frac{\partial^{3}u}{\partial x^{3}}$ (S4)
+$\frac{\partial u}{\partial t}=-u\frac{\partial u}{\partial x}-\frac{\partial^{3}u}{\partial x^{3}}$ (9)
 We got the 8 sets of data: 5 for training, 1 for validation, and 2 for testing. The data sets had spatial domain size $x\in[0,64]$, where $\Delta t$ is 15 times $\delta t$ and $\delta t=0.01s$.
@@
 The 2D Burgers’ equation is commonly employed as a benchmark model for comparing and evaluating different computational algorithms, and describes the complex interaction between nonlinear convection and diffusion processes in the way like:
-$\frac{\partial u}{\partial t}$ $=-uu_{x}-vu_{y}+\nu\nabla^{2}u,$ (S5)
-$\frac{\partial v}{\partial t}$ $=-uv_{x}-vv_{y}+\nu\nabla^{2}v.$ (S6)
+$\frac{\partial u}{\partial t}$ $=-uu_{x}-vu_{y}+\nu\nabla^{2}u,$ (10)
+$\frac{\partial v}{\partial t}$ $=-uv_{x}-vv_{y}+\nu\nabla^{2}v.$ (11)
 The $u$ and $v$ are the fluid velocities and $\nu$ denotes the viscosity coefficient. In this case, we choose $\nu=0.005$, where $\delta t=0.001s$.
 2D FitzHugh-Nagumo Equation. The FitzHugh-Nagumo system can be described by the equation:
-$\frac{\partial u}{\partial t}$ $=\mu_{u}\nabla^{2}u+u-u^{3}-v+\alpha,$ (S7)
-$\frac{\partial v}{\partial t}$ $=\mu_{v}\nabla^{2}v+(u-v)\beta.$ (S8)
+$\frac{\partial u}{\partial t}$ $=\mu_{u}\nabla^{2}u+u-u^{3}-v+\alpha,$ (12)
+$\frac{\partial v}{\partial t}$ $=\mu_{v}\nabla^{2}v+(u-v)\beta.$ (13)
 The coefficients $\alpha=0.01$ and $\beta=0.25$, governing the reaction process, take distinct values, while the diffusion coefficients are $\mu_{u}=1$ and $\mu_{v}=100$.
@@
 Gray-Scott Equation. The Gray-Scott equations describe the temporal and spatial variations of chemical concentrations in reaction-diffusion systems, which can be described by the equation:
-$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+F(1-u),$ (S9)
-$\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(F+k)v.$ (S10)
+$\frac{\partial u}{\partial t}$ $=D_{u}\nabla^{2}u-uv^{2}+F(1-u),$ (14)
+$\frac{\partial v}{\partial t}$ $=D_{v}\nabla^{2}v+uv^{2}-(F+k)v.$ (15)
 Here, in the two-dimensional case, $D_{u}$ and $D_{v}$ represent the diffusion coefficients of the two substances, with specific values of $D_{u}=2.0\times 10^{-5}$ and $D_{v}=5.0\times 10^{-6}$.
@@
 In the 2D Gray-Scott case, we got 5 trajectories for training, 1 trajectory for validation and 3 trajectories for testing, where $\delta t=0.5s$.
-In the three-dimensional case, we have the parameters: $DA=0.2$, $DB=0.1$, $F=0.025$, and $k=0.055$.
+In the three-dimensional case, we have the parameters: $D_{u}=0.2$, $D_{v}=0.1$, $F=0.025$, and $k=0.055$.
 We got 3 trajectories for training, 1 trajectory for validation and 2 trajectories for testing, where $\Delta t=15\delta t$ and $\delta t=0.25s$.
-Appendix H Supplement results
+G: Supplement results
+Models RMSE MAE
+PIMRL 0.0050 0.0039
+PeRCNN 0.0137 0.0109
+FNO 0.0626 0.0502
+FNO-coarse 0.0415 0.0332
+Table 7: The experiments between PIMRL and baseline models on the Burgers example of FNO benchmark.
 Considering the constraints on the length of the main text, we have placed some important supplementary results in the appendix of the article.
-First, we conducted additional experiments on the FN equation, expanding our tests to include more advanced baselines such as F-FNO, U-NO, and MWT (shown in table S2).
+First, we conducted additional experiments on the FN equation, expanding our tests to include more advanced baselines such as F-FNO, U-NO, and MWT (shown in table 4).
 The experimental results indicate that data-driven methods do not perform well on small training datasets, which further corroborates the effectiveness of the PIMRL proposed in our paper.
+Figure 9: Computational time for comparison
+Inference Time
+As shown in Figure 9, PIMRL significantly reduces the computational time cost.
+Compared to traditional methods such as Direct Numerical Simulation (DNS), our framework is substantially faster, demonstrating a significant improvement in computational efficiency.
+Additionally, PIMRL not only delivers superior predictive accuracy, but also outperforms PeRCNN in terms of computational efficiency.
+The numerical methods used for DNS are consistent with the parameter settings during data generation.
+The difference lies in the required prediction time length, which needs to be the same as that of PIMRL and PeRCNN for a fair comparison.
+Overall, PIMRL achieves fast and accurate long-term predictions, making it highly advantageous for both precision and real-time applications.
 To demonstrate the stability of PIMRL, we have included corresponding error bars in the appendix.
@@
 This analysis offers insights into the computational efficiency and resource requirements of our model relative to existing approaches, further highlighting the practical advantages of PIMRL in terms of performance optimization and scalability.
-Table S2: The Result of U-NO, F-FNO, MWT and FNO in the FN Case.
-Metrics U-NO F-FNO MWT FNO
-RMSE 0.3675 0.2280 0.3494 0.1878
-MAE 0.1465 0.1350 0.2228 0.1634
-Table S3: Results with Error Bar under RMSE metric on all Cases.
 Model PIMRL PeRCNN
-Kdv $0.0457\pm 0.0053$ $0.0942\pm 0.0082$
-Burgers $0.0068\pm 0.0006$ $0.0075\pm 0.0008$
-2DGS $0.0133\pm 2.4\times 10^{-12}$ $0.0455\pm 1.9\times 10^{-11}$
+Kdv $0457\pm 0.0053 0.0$$0942\pm 0.0082$
+Burgers $0068\pm 0.0006 0.0$$0075\pm 0.0008$
+2DGS $0133\pm 2.4\times 10^{-12} 0.0$$0455\pm 1.9\times 10^{-11}$
 FN $0.1349\pm 0.0040$ $0.1591\pm 0.0061$
-3DGS $0.0381\pm 0.0015$ $0.0532\pm 0.0027$
-Table S4: The experiments between PIMRL and baseline models on the Burgers example of FNO benchmark.
-Models RMSE MAE
-PIMRL 0.0050 0.0039
-PeRCNN 0.0137 0.0109
-FNO 0.0626 0.0502
-FNO-coarse 0.0415 0.0332
-Table S5: Running Time, Parameter Size, and GPU Memory for PIMRL, U-NO, MWT, FNO, and ConvLSTM
-Model Running Time Parameter Size GPU Memory
-PIMRL 9 s 3.33 M 1728 M
-U-NO 7 s 15.29 M 2320 M
-MWT 12 s 0.09 M 1732 M
-FNO 2 s 8.39 M 1580 M
-ConvLSTM 5 s 3.32 M 1708 M
+3DGS $0381\pm 0.0015 0.0$$0532\pm 0.0027$
+Table 8: Results with Error Bar under RMSE metric.
+H: Limitations and Future Work
+While our proposed PIMRL framework demonstrates significant promise, we acknowledge several limitations that pave the way for future research.
+Firstly, the performance sensitivity of the framework to the duration of the data gap between burst-sampling intervals has not been exhaustively studied.
+Understanding the critical point at which the macro-module’s predictive error becomes too large for the micro-module to correct is essential for practical applications.
+Secondly, our current experiments were conducted on clean, numerically generated data.
+The framework’s robustness to varying levels of noise, which is ubiquitous in real-world sensor measurements, remains an open question for investigation.
+Looking ahead, our efforts will address these limitations and extend the framework’s capabilities.
+We plan to explore more advanced architectures for the macro-scale module to enhance both computational efficiency and predictive accuracy, particularly for highly complex dynamical systems.
+Furthermore, we will work to integrate super-resolution techniques, enabling the model to generate high-fidelity predictions on finer grids than those it was trained on.
+Systematically evaluating the model’s performance under noisy and more sparsely sampled conditions will also be a priority.
+These advancements will be crucial for unlocking PIMRL’s full potential in modeling intricate physical phenomena and broadening its impact across scientific computing.